### v1

# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

Abstract
Large language models (LLMs) have become pervasive in our everyday life. Yet, a fundamental obstacle prevents their use in many critical applications: their propensity to generate fluent, human-quality content that is not grounded in reality. The detection of such hallucinations is thus of the highest importance. In this work, we propose a new method to flag hallucinated content, MMD-Flagger. It relies on Maximum Mean Discrepancy (MMD), a non-parametric distance between distributions. On a high-level perspective, MMD-Flagger tracks the MMD between the generated documents and documents generated with various temperature parameters. We show empirically that inspecting the shape of this trajectory is sufficient to detect most hallucinations. This novel method is benchmarked on two machine translation datasets, on which it outperforms natural competitors.
1 Introduction
Neural probabilistic modeling (bengio_et_al_2003, ) in an auto-regressive framework has become the de facto paradigm for natural language processing. Powered by the transformer architecture (Vaswani-2017-attention, ), such neural networks have achieved impressive performances in text translation (kocmi_et_al_2024, ), abstractive text summarization (shakil_farooq_kalita_2024, ), generative question-answering (daull_et_al_2023, ), and many other applications. More recently, the trend has been to bring forward so-called Large Language Models (LLMs) such as GPT (openai_2024, ), which can tackle simultaneously nearly all these different tasks. Popularized by ChatGPT (a chatbot version of GPT), they are used daily by dozens of millions of users (Yuyang-2023-BigData, ).
However, LLMs have a key limitation which prevents their use in critical applications: their propensity to generate hallucinations. Commonly characterized by the generation of fluent but factually inaccurate or unsubstantiated content, these are observed in all models (ji_et_al_2023, ), although rare (Guerreiro-2023-looking, ). Within the domain of machine translation, the focus of this work, this phenomenon manifests as translations that, while grammatically correct, deviate semantically from the contextual information present in the source language (Guerreiro-2023-looking, ). We provide an example in Table 1. Hallucinated content is a critical drawback, especially in domains where information accuracy is essential, such as law Benkirane-2024-machine . In these fields, getting wrong (yet seemingly correct) information even once can undermine users’ trust permanently. Detecting hallucinations at test time remains a challenging task.
Focusing on standard encoder-decoder architecture, it is believed that the decoding step is one of the main causes of hallucinations (ji_et_al_2023, ). In practice, beam search is commonly used to generate the output token sequence, given its superior performance, even with respect to maximum a posteriori (Meister-etal-2020-beam, ). This is particularly true for machine translation tasks (Guerreiro-2023-looking, ; Xu-etal-2023-understanding, ). Nevertheless, beam search tends to suffer from the so-called likelihood trap (das-balke-2022-quantifying, ): a high-likelihood token sequence often results in a low-quality token sequence.
An alternative approach is simply to sample the next token according to the probability distribution given by the decoder’s softmax layer at each step. The main advantage in doing so is the stochasticity of outputs, giving rise to many diverse output documents from a single input. A further refinement is to introduce a temperature parameter ($\tau$ in our notation), which controls the randomness of the next token prediction; a low value (say $<0.1$) corresponding roughly to a probability distribution concentrated on the most likely token, and a higher value (say $>1$) making the distribution near uniform. Thus, it is not surprising that, in practice, low temperatures are reported to induce low variability in the sampled token sequences (Peeperkorn-etal-2024, ), while high temperatures result increase in the hallucination rate (Renze-2024-effect, ; Huang-2025, ).
Table 1: Examples of hallucinated translation from the LFAN-HALL dataset generated by beam search ($y_{\text{beam}}$), and stochastic samples with varying temperature parameters ($y_{\text{sto}}^{\tau}$). The MMD distances between $y_{\text{beam}}$ and $y_{\text{sto}}^{\tau=0.5}$ as a function of $\tau$ can be visualized as the red line in Figure 1.
Source (German) Gutes Angebot bei Techno, House und Minimal Techno.
Reference translation Good choice of Techno, House and Minimal Techno tracks.
$y_{\rm beam}$ The staff were very friendly and helpful. The room was clean and comfortable.
$y^{\tau=0.1}_{\rm sto}$ The hotel is very close to the city centre and the train station.
The hotel is very well located, close to the beach and the beach.
$y^{\tau=0.5}_{\rm sto}$ The staff was very helpful and friendly.
The hotel is very close to the city centre and the metro station.
$y^{\tau=0.8}_{\rm sto}$ It had the best overall prices. It was close to the beach and business.
The breakfast buffet was very good häufude and the location is ideal for getting to the town.
The central objective of this study is to compare the default output of a language model (generated using beam search) with stochastic samples obtained under varying temperature. At low temperature settings, we observe that the sampled outputs are highly similar to the default output if the model is not hallucinating, and moderately dissimilar if it is. We believe this occurs because of the following reason: when a clear optimal candidate token exists, both sampling (at low temperatures) and beam search tend to favor the most probable token. However, in cases where no single candidate clearly dominates (often indicative of hallucination), the outputs are more likely to diverge. In contrast, at high temperature, generation becomes increasingly random, and sampled outputs tend to be substantially different from the default output regardless of the model’s behavior. This pattern is illustrated in Table 1. As a consequence, plotting the similarity between the default output and the stochastic outputs yields two distinct curves: either a monotonously increasing trajectory in the non-hallucinating scenario, or a $U$-shape curve in the hallucinating scenario. This can be visualized in Figure 1.
Drawing on these findings, we introduce a novel hallucination detection method: MMD-Flagger. The key idea of the method is to simply i) plot the similarity between default output and stochastic samples as a function of the temperature, and ii) automatically detect whether this curve is monotonously increasing or $U$-shaped. In the latter case, we flag the output as hallucination. As a measure of similarity, we propose to use the Maximum Mean Discrepancy (MMD) (Gretton-2012a, ). As a shape detection mechanism, we simply estimate the inflection point of the trajectory and check whether it is located near 00 (see Figure 1).
The paper is organized as follows. In Section 2, we detail the operating procedure of MMD-Flagger after recalling the necessary background on temperature sampling and MMD. We benchmark our method against several natural competitors in Section 3, before concluding in Section 5. The codebase is available at https://github.com/Kensuke-Mitsuzawa/MMD-Flagger.
Figure 1: MMD-Flagger at work on the LFAN-HALL dataset: two outputs, generating two MMD trajectories. In red, the trajectory for the hallucinated translation (see Table 1), in green, the trajectory corresponding to the correct one. Text boxes show examples of $y_{\rm beam}$ and samples from $Y_{\rm sto}^{\tau}$ (two of 25 samples), generated by stochastic decoding with temperatures $\tau$. We flag $y_{\rm beam}$ as a hallucination when the minimum value of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ (the vertical black dashed line) falls within the red region. Otherwise, the output is deemed plausible (minimum falling into the green region).
1.1 Related work
Hallucination detection is a very active field, with a wide range of approaches proposed in recent years. Let us review the most related work; We refer to (ji_et_al_2023, ) for a more complete survey of mitigation methods in various domains.
Uncertainty quantification. Following the taxonomy outlined by (Huang-2025, ), MMD-Flagger falls under the umbrella of uncertainty estimation, which aims to quantify a model’s confidence in its output. Stemming from the observation that models are more confident when they do not hallucinate, the Monte-Carlo dropout-based MC-Dsim (Fomicheva-2020-unsupervised, ; Zerva-2021-unbabel, ) estimates the stability of generated token sequences. A similar approach is to inspect the log-probability of output sequences (Guerreiro-2023-looking, ) and to flag as hallucination if this log-probability is too low (Seq-Logprob). Alternatively, quality estimation methods (Rei-etal-2022-cometkiwi, ; Feng-etal-2022-language, ) correlate low-quality translations with hallucinations, though they are not originally designed for hallucination detection. Empirical evidence also suggests that combining multiple detectors improves detection accuracy (Guerreiro-2023-looking, ; Himmi-2024-enhanced, ). In the context of QA, (Varshney-2023stitchtimesavesnine, ) proposes to use the minimum token probability as an indicator of uncertainty, based on the premise that lower token probabilities suggest higher uncertainty. We compare MMD-Flagger to MC-Dsim and Seq-Logprob in Section 3.
External knowledge. Beyond uncertainty estimation, several approaches aim to detect hallucination by leveraging external knowledge. For example, (Min-etal-2023-factscore, ) introduce factuality scores derived from knowledge graphs. In machine translation, (Xu-etal-2023-understanding, ) propose a supervised hallucination detection model trained on synthetic pseudo-hallucination datasets. The LMvLM framework (Cohen-etal-2023-lm, ) employs an examiner language model to interrogate an examinee model through multi-turn interactions, identifying potential inconsistencies. Similarly, the Chain-of-Verification method (dhuliawala-etal-2024-chain, ) makes the model’s reasoning process explicit, enabling humans to assess the validity of its outputs. We do not compare against such methods since MMD-Flagger does not use external information.
Model internals. Building on the assumption that models “know when they hallucinate” (duan_et_al_2024, ), other approaches look at models’ internal states to detect hallucinations. (Ferrando-etal-2022-towards, ; Dale-2022-detecting, ) measure how much the input contributes to the output and show that one can detect hallucination when the source sentence contribution is low. It is also possible to analyze input-output relationships from an approach using optimal transport (Guerreiro-etal-2023-optimal, ). We showcase in Appendix G how MMD-Flagger can be adapted to leverage information from the model internals. However, the more consistent experimental results are achieved using only the model’s outputs.
Kernel methods and LLMs. Another key ingredient of our approach is the use of MMD to compute similarities between outputs of LLMs. To the best of our knowledge, only one prior study connects MMD with LLMs. (Zhangs-2024MMDMP, ) propose a two-sample testing framework that employs MMD as the test statistic to distinguish between human-written and machine-generated texts. Thus, their goal is not hallucination detection, but rather authenticity classification, making the application of MMD to hallucination detection a novel contribution of our work. More similar to our approach, (Nikitin-2024-kernel, ) proposes kernel language entropy, a kernel-based method to quantify the uncertainty in LLM outcomes, and hallucination detection is proposed based on the entropy score (Farquhar-2024detecting, ).
2 Methodology
In this section, we describe the proposed method, MMD-Flagger. Given an input document $x$, the starting point of the method is to generate the output $y=f(x)$ to be inspected, using beam search. The next step is to generate additional documents by stochastic sampling at various temperature parameters, which is detailed in Section 2.1. Then, the MMD between $y$ and the newly created groups is computed for each temperature parameter. We detail this computation in Section 2.2. At this point, for a given input / output pair $(x,y)$, we have obtained a series of MMD values corresponding to different temperature settings. We explain our procedure to inspect this trajectory in Section 2.3.
2.1 Stochastic sampling and temperature
In this section, we present the encoder-decoder architecture considered in this work and detail the role of the temperature in stochastic sampling.
Encoder-decoder architecture.
We consider the encoder-decoder architecture (Sutskever-2014, ) with self-attention mechanism (Vaswani-2017-attention, ). Given an input token sequence $x$, the encoder $e$, which consists in multiple layers of self-attention and feed-forward networks encodes the sequence into a representation with fixed length $T$. The decoder then generates the output sequence autoregressively, conditioned on both the encoded fixed-length vector and the already processed tokens. More precisely, each output token is generated as
$Y_{t}=\Phi(Y_{t-1},\ldots,Y_{1},e(x))\,,$
where $\Phi$ is a (potentially) stochastic sampler. This process iterates until the EOS token is sampled. In our framework, the way sampling operates is based on the assumption that the decoder is equipped with a softmax layer converting logits scores $\mathbb{R}^{T\times V}$ into a discrete probability distribution $P\in\mathbb{R}^{V}$, where $V$ is the size of the dictionary.
Argmax. Arguably, the most natural thing to do is to select the token associated with the highest $P_{i}$ as output. We refer to this as argmax. Simply picking the most likely token comes with a lot of rigidity, since some tokens can attract a lot of mass, such as the EOS token. Alternatively, it is very common to sample a random token using this distribution. There are many ways to do so.
Stochastic sampling. A first possibility is simply to sample according to $P$. That is, sample a random variable with distribution given by $P$ and select the associated token. We denote $Y_{\text{sto}}=\text{LLM}_{\text{sto}}(x)$ the random output obtained by stochastic sampling.
Temperature parameter. A variation on the previous theme is to introduce a temperature parameter $\tau>0$. The softmax transformation is modified into
$P_{i}=\frac{\mathrm{exp}(z_{i}/\tau)}{\sum_{j=1}^{V}\mathrm{exp}(z_{j}/\tau)}\,,$ (1)
where $z_{i}$ represents the logit score for the $i$-th token. As recalled in Section 1, high temperature gives a lot of variability (Yehuda-etal-2024, ) (equivalent to uniform sampling in the limit), while low temperature degrades to argmax sampling.
Beam Search. Rather than considering only the most likely next token, beam search allows for a broader search over possible sequences by keeping multiple candidates. This principle underlies beam search, a deterministic decoding algorithm widely used in the literature (Meister-etal-2020-beam, ). Formally,
$\begin{split}S_{t}&=\operatorname{argmax}_{S^{\prime}\subseteq\mathcal{B}_{t},|S^{\prime}|=w}\log p_{\theta}(S^{\prime}|x),\\ \text{where}\qquad\mathcal{B}_{t}&=\{s_{t-1}\circ s\mid s\in\mathcal{V}\text{ and }s_{t-1}\in S_{t-1}\}\,.\end{split}$
In the previous display, $S$ is the stack of candidate token sequences maintained during beam search, $S_{t}$ is the stack at the $t$-th time step, $\mathcal{B}_{t}$ is the candidate set of next possible token sequences, and $\mathcal{V}$ indicates the vocabulary list, and $s_{t-1}\circ s$ is the operation of adding a sampled token $s$. Beam search usually starts with $S_{0}=\{\text{BOS}\}$, and holds $w$-candidates of token sequences in $S$. Beam search can keep the $w$-best path of tokens in selecting the $t$-th token, while the argmax sampling simply chooses the token with the highest probability value. We denote by $y_{\text{beam}}=\text{LLM}_{\text{beam}}(x)$ the token sequence generated by beam search. Note that it is also possible to use beam-search with a temperature parameter, but we do not do so in this work, as is typical. While aiming to find the most likely token sequence, beam search is known to sometimes generate low-quality outputs having a high likelihood (das-balke-2022-quantifying, ). Stochastic sampling offers a potential mitigation for this issue by introducing diversity into the generation process. We aim to compare the sequences generated by beam search with those produced through stochastic sampling. For a temperature parameter $\tau\in T$, we collect a set of token sequences $Y^{\tau}_{\rm sto}=\{\text{LLM}^{\tau}_{\text{sto}}(x)\}_{n=1}^{N}$ generated by stochastic sampling. To summarize, in this work we compare $y_{\text{beam}}$ and $Y^{\tau}_{\rm sto}$ for various values of $\tau$.
Alternative sampling strategies. To conclude this section, we note that there exist many other alternative sampling strategies, such as top-k (fan-etal-2018-hierarchical, ) or nucleus sampling (Holtzman-2020, ). They can mitigate the hallucinations by placing constraints on sampling the next tokens (Huang-2025, ).
2.2 Maximum Mean Discrepancy (MMD)
Figure 2: Illustration of the maximum mean discrepancy (MMD). MMD measures the distance between two probability distributions by mapping them into a Reproducing Kernel Hilbert Space (RKHS) and computing the distance.
Maximum Mean Discrepancy (MMD) is a measure of distance between two probability distributions. It relies on a positive semi-definite kernel (Hofmann-2008, ; Smola-2009, ) to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points (Muandet-2017, ). A large MMD value indicates significant differences between two distributions. We depict an intuitive presentation of the MMD in Figure 2.
Formal definition. Formally, for a given kernel $k$, the MMD between two distributions $P$ and $Q$ is defined as
$\begin{split}{\rm MMD}_{k}^{2}(P,Q):=\ \mathbb{E}_{A,A^{\prime}\sim P}[k(A,A^{\prime})]+\mathbb{E}_{B,B^{\prime}\sim Q}[k(B,B^{\prime})]-2\mathbb{E}_{A\sim P,\ B\sim Q}[k(A,B)]\,,\end{split}$ (2)
where $A,A^{\prime}\in\mathbb{R}^{D}$ are independent random vectors drawn from $P$, and $B,B^{\prime}\in\mathbb{R}^{D}$ are drawn from $Q$.
Empirical estimate. Given i.i.d. samples ${\bf A}=\{A^{1},\dots,A^{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and ${\bf B}=\{B^{1},\dots,B^{m}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, one can estimate the MMD by replacing the expectations in Eq. (2) by the corresponding empirical averages:
$\begin{split}&\widehat{\rm MMD}^{2}_{U}({\bf A},{\bf B}):=\frac{1}{n(n-1)}\sum_{1\leq i\neq i^{\prime}\leq n}k(A^{i},A^{i^{\prime}})\\ &+\frac{1}{m(m-1)}\sum_{1\leq j\neq j^{\prime}\leq m}k(B^{j},B^{j^{\prime}})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(A^{i},B^{j}).\end{split}$ (3)
This is an unbiased estimate of the MMD Eq. (2), and converges at the rate $\mathcal{O}_{\mathbf{P}}(\min(n,m)^{-1/2})$ as $n,m\to\infty$ provided that $\sup_{x\in\mathbb{R}^{d}}k(x,x)<\infty$ (Gretton-2012a, , Theorem 10).
MMD in our setting. Recall that our goal is to measure the similarity between the outputs of $\text{LLM}_{\rm beam}$ and those of $\text{LLM}_{\rm sto}^{\tau}$. To do this, we consider $\text{LLM}_{\rm beam}$ and $\text{LLM}_{\rm sto}^{\tau}$ as two probability distributions, and use Eq. (3) to get an estimate of the MMD. Note that, since $\text{LLM}_{\rm beam}$ deterministically generates a sequence, it is simply a constant random variable.
Vector representation of outputs.
Since most kernels (especially the ones used in this work) take vector values as input, we need to convert outputs into vector representation. The most natural way is to resort to the word embedding matrix $E\in\mathbb{R}^{V\times D}$ used by our encoder-decoder. Thus each token sequence $y$ is associated with a sequence of embedding vectors
$(e_{1},\ldots,e_{T})=E_{y_{1},:},\ldots,E_{y_{T},:}\in\mathbb{R}^{T\times D}\,.$
We introduce two aggregations of $T$-length vector representations and convert them into a fixed-length vector $h$; avg and concat. The avg aggregation is computing the average over $T$ vectors, namely
$h=\frac{1}{T}\sum_{t=1}^{T}e_{t}\in\mathbb{R}^{D}\,.$
The other possibility we investigate is the concat aggregation joins the $T$ embeddings to form
$h=[e_{1};e_{2};\ldots;e_{T};\underbrace{e_{\rm pad};e_{\rm pad};\ldots;e_{\rm pad}}_{T_{\max}-T}]\in\mathbb{R}^{T_{\max}\cdot D}\,,$
where $e_{\rm pad}=\mathbf{0}_{D}$, and $T_{\rm max}$ is given as a hyperparameter. Having the fixed-length vector representation, we denote the set of fixed-length vectors as $H:=\{h_{1},\ldots,h_{n}\}$.
Kernel choice. In this work, we focus on the linear and the Gaussian kernel. The linear kernel (Dot in the following) $k^{\text{lin}}(a,b)=a^{\top}b$ is the simplest choice, and gives rise to simple measures of similarity. Indeed, it is easy to show that the MMD in that case corresponds to the $L^{2}$ distance between $y_{\text{beam}}$ and the barycenter of $Y^{\tau}_{\rm sto}$. On the other side, the Gaussian kernel is often preferred in machine learning applications, notably because of the flexibility of its parameter tuning. For any $a,b\in\mathbb{R}^{D}$, it is defined as
$k^{\text{Gau}}(a,b)=\exp\left(\frac{-\|a-b\|^{2}}{2\gamma^{2}}\right)\,.$ (4)
In the previous display, the bandwidth parameter $\gamma>0$ controls the scaling of input distances before taking the exponential. The median heuristic (Garreau2018-largesampleanalysismedian, ), which sets the length scale to the median of pairwise distances, is commonly used in kernel methods due to its empirical effectiveness. We explored various settings of length scale selections, and selected the best percentile at $25$-th percentile, which is reported in Appendix F.
To conclude this section, we note that MMD is a good choice when the data dimensionality is high and the available sample size is limited (Gretton-2012a, ). These MMD’s advantages are well-suited to our setting since the vectorized outputs can have quite a high dimension (especially in the concat setting), while the number of sampled sequences $H^{\tau}_{\text{sto}}$ is strongly limited by our compute budget.
2.3 MMD Flagger
Now, we are equipped with all the machinery required to introduce our proposed method. MMD-Flagger is a simple rule-based approach checking MMD distances $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ for a grid of temperature parameters $T=\{\tau_{1},\dots,\tau\}$. By default, we propose to take $T=\{0.1,\ldots,1.0\}$ with a step of $0.1$. For each $\tau\in T$, we query the model to sample token sequences $Y^{\tau}_{\rm sto}:=\{{y^{\tau}_{\rm sto}}\}_{n=1}^{N}$ and its vector representations $H^{\tau}_{\rm sto}:=\{h_{\rm sto}^{\tau}\}_{n=1}^{N}$ by the stochastic sampling with temperature parameter $\tau$, as described in Section 2.1. We then compute the MMD distances defined by Eq. (3). Namely,
$\mathcal{T}_{\widehat{MMD}}:=(\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto}),\ldots,\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto}))\,.$
Since an empirical MMD estimate (Eq 3) converges to the true MMD as the sample size increases, the variability in the computed MMD trajectory gradually disappears, as empirically shown in Appendix H.
After computing these values, we detect the inflection point of this trajectory and flag it as a hallucination if it is close to 00. To detect this inflection point, we proceed in the following way: first, we consider a minimal temperature $\tau_{0}$, set by default at $0.11$. Then we compute $\tau_{\text{min}}$ the temperature parameter realizing the minimal MMD distance between the reference and the stochastic samples. In other words, $\tau_{\min}$ is the index of $\mathcal{T}_{\widehat{MMD}}$ with the minimal value. Finally, we flag as hallucination if $\tau_{\min}$ is greater than or equal to $\tau_{0}$. Figure 1 showcases an example of the method at work in both the hallucination and the non-hallucination scenario. We collect the operating procedure in Algorithm 1.
Input: LLM, source input $x$, temperature parameters $T$, number of samples $N$, a minimum temperature for detecting the hallucination $\tau_{0}$
Obtain beam-search translation $h_{\text{beam}}=\text{LLM}_{\text{beam}}(x)$
foreach $\tau\in T$ do
       Repeat stochastic translation in $N$ times and obtain $H_{\text{sto}}^{\tau}=\text{LLM}_{\text{sto}}^{\tau}(x)$
       Compute the MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ using Eq. (3)
end foreach
$\tau_{\min}=\arg\min(\mathcal{T}_{\widehat{MMD}})$
if $\tau_{\min}>\tau_{0}$ then
       return hallucination
Algorithm 1 MMD-Flagger
3 Empirical Assessment
In this section, we evaluate MMD-Flagger on two machine translation datasets. All experiments were done on a single NVIDIA Tesla V100 GPU. We report more details on the experimental setup in Appendix B.
Table 2: Hallucination detection performance reported as “Recall (Precision).” We mark the best recall in bold. MMD-Flagger matches existing baseline in terms of recall on the LFAN-HALL dataset, and outperforms in terms of recall and precision on the Halomi dataset. We refer to Section 3.3 for more details.
LFAN-HALL Halomi
TNG 0.069 (1.000) 0.002 (1.000)
Baselines MC-Dsim 0.815 (0.182) 0.553 (0.429)
Seq-Logprob 0.850 (0.189) 0.381 (0.273)
MMD-Flagger (avg-Gaussian) 0.816 (0.129) 0.825 (0.281)
Ours MMD-Flagger (avg-Dot) 0.803 (0.124) 0.779 (0.278)
MMD-Flagger (concat-Gaussian) 0.850 (0.134) 0.824 (0.278)
MMD-Flagger (concat-Dot) 0.833 (0.130) 0.817 (0.282)
3.1 Datasets and metrics
LFAN-HALL (Guerreiro-2023-looking, ) is a hallucination detection benchmark dataset for the machine translation task from German to English. The dataset consists in $3,413$ machine translations obtained by a transformer-based encoder-decoder model relying on the fairseq package Ott-2019-fairseq and beam search, together with associated reference translation and human-annotated hallucination labels. Since we observed some inaccurate annotation labels, we undertook a review of the annotated labels and made corrections. The relabeling process is detailed in Appendix A. The dataset defines three types of hallucination labels occupying approximately $8\%$ of all translations, and we unify these three labels as “hallucination.” In this dataset, the beam search width is fixed to $5$.
Halomi (Dale-2022-detecting, ) is a hallucination detection benchmark dataset consisting over $3,000$ sentences in $9$ languages. Similarly to LFAN-HALL, Halomi consists of beam search translations manually annotated for hallucination. The dataset defines three types of hallucinations that occupy approximately $26\%$ of the dataset, and we again unify them as “hallucination.” Following the setting of (Himmi-2024-enhanced, ), we use a distilled version of the multilingual translation model (Costa-2022-no, ), totaling 600M parameters. We use the open source Hugging Face implementation. Beam search width was also set to $5$.
Evaluation Metrics. We use recall and precision as evaluation metrics. They are computed based on the binary labels associated with hallucination / not hallucination given by the datasets. We treat the “hallucination” label as true label. Recall is defined as $\frac{T_{p}}{T_{p}+F_{n}}$, and precision is defined as $\frac{T_{p}}{T_{p}+F_{p}},$ where $T_{p}$ denotes the number of true positives, $F_{p}$ the number of false positives, and $F_{n}$ the number of false negatives.
3.2 MMD-Flagger settings
As temperature parameters, we set $T=\{0.1,0.2,\ldots,0.9,1.0\}$. Previous work (Peeperkorn-etal-2024, ) used a smaller value $\tau<0.1$, however, in our setting, the minimum $\tau$ is set to $\tau=0.1$ since the fairseq package runs into numerical problems whenever $\tau<0.1$. The set of stochastic sampling sequences $Y_{\text{sto}}^{\tau}$ and corresponding vector representations $H_{\rm sto}^{\tau}$ consist of $25$ examples for each temperature parameter. We obtain the embedding representation from the word embedding layer of the given LLM. The embedding dimensions are $512$ for LFAN-HALL and $1,024$ for Halomi.
As described in Section 2.2, we considered the linear(Dot) and Gaussian (Gaussian) kernels in our experiments. To pick the bandwidth $\gamma$ for the Gaussian kernel, we randomly select $200$ (correct) translations as calibration data. The $T_{\rm max}$ of concat is set to the maximum token length among $200$ translations. We conduct experiments with various configurations of choosing the length scale and select the $25$-th percentile of $L^{2}$-distance, instead of the usual median heuristic (Garreau2018-largesampleanalysismedian, ). Supporting experiments for this choice can be found in Appendix F. For the Halomi dataset, selecting the calibration data and calculating the length scale is done per source and target language pair.
We also benchmarked MMD-Flagger on internal representations of the model (namely, taking as $h$ vectors of the intermediate hidden layers). We did not find a consistent pattern in our results, while some internal representations yield better experimental results, see Appendix G.
3.3 Baseline Methods
We considered the baselines proposed by (Guerreiro-2023-looking, ); one rule-based method (TNG) and two uncertainty estimation methods (MC-DSim, Seq-Logprob). Let us describe them briefly.
Top n-gram count (TNG(Raunak-etal-2021-curious, ) is a rule-based method that compares frequencies of $n$-grams between the source and the translation text. TNG flags the translation as the hallucination if certain $n$-gram counts have significant differences. TNG is designed to detect the oscillation type of hallucination; thus, it struggles to flag strong or fully-detached hallucination cases. We set $n=4$ following (Guerreiro-2023-looking, ).
Dissimilarity of Monte-Carlo hypotheses (MC-DSim(Fomicheva-2020-unsupervised, ; Zerva-2021-unbabel, ) is the average similarity scores between $y_{\text{beam}}$ and $Y_{\rm MC}$, defined by $\frac{1}{N}\sum_{i=1}^{N}\text{SIM}(y_{\text{beam}},Y_{\rm MC}),$ where $Y_{\rm MC}$ is a token sequence generated by the beam search using Monte-Carlo Dropout (Gal-2016-dropout, ), and $N$ is the number of Monte-Carlo Dropout generations. Following (Guerreiro-2023-looking, ), we set $N=10$ and use the METEOR metric (Banerjee-2005-meteor, ) of the nltk package (Bird-loper-2004-nltk, ) to compute the similarity score.
Sequence log-probability (Seq-Logprob) is the average of log-probability, defined by $\frac{1}{T}\sum_{k=1}^{T}\log\mathbb{P}\left(y_{k}\middle|y_{<k},x,\theta\right),$ where $x$ is the source text, $y_{k}$ is the $k$-th token in the translation using the beam search, $y_{<k}$ is the previous tokens, $T$ is the length of tokens in the translation, and $\theta$ is the model parameters of a LLM.
MC-DSim and Seq-Logprob are both using a fixed threshold to detect hallucination. As Guerreiro-2023-looking suggests, we fix the threshold at the $40$-th percentile among a given dataset. On the LFAN-HALL dataset, the threshold selection is performed using the whole dataset, while we compute the threshold per language pair of the source and target language on the Halomi dataset. The threshold selection is described in more depth in Appendix C.2.
3.4 Discussion
Experimental results. Table 2 lists the precision and recall scores on the benchmark datasets. MMD-Flagger (concat-Gaussian) matches the best baseline in terms of recall on the LFAN-HALL dataset (Seq-Logprob), and outperforms it (again Seq-Logprob) in terms of recall and precision on the Halomi dataset. We see these results as very promising, and observe that they are robust with respect to changes in the kernel and aggregation method.
Concatenation versus averaging. When comparing avg and concat, the concat aggregation generally yielded slightly superior or comparable performance to avg. This trend is plausible, as concat retains all individual embeddings, whereas avg aggregates information through a mean operation, potentially losing finer-grained details. Indeed, concat relies on a much larger vector representation of the output documents, of size $D\cdot T_{\text{max}}$. This can become an issue for longer outputs or if the dimension of the word embedding is too high. For instance, in our experiments with the LFAN-HALL dataset, the selected maximum token length ($T_{\text{max}}$) by the calibration data is $105$ tokens. Given that the dimensionality of a word embedding vector $D=512$, the resulting concat vectors have dimension $53,760$. In contrast, the avg aggregation maintains a fixed-length vector size regardless of the token length. Thus, the avg aggregation is generally more practical, even though detection scores can be slightly lower than those of the concat aggregation.
4 Discussion
Limitations. The first limitation of MMD-Flagger lies in its reliance on the source of hallucination. The method is designed to detect hallucinations by identifying discrepancies between beam search and stochastic sampling outputs, as reflected in the trajectory of MMD distances. Consequently, if the sampled sequences $Y_{\text{sto}}^{\tau}$ are highly similar to the beam search output $y_{\text{beam}}$ for low temperatures, hallucinations can go undetected. It may be the case for an irreducible proportion of hallucinations.
Another limitation is the computational cost of MMD-Flagger, as it requires repeated stochastic sampling across multiple temperature settings. This cost scales with the size of LLM and the length of the generated sequence, making the method less suitable for very large models or long outputs. The computational cost of the MMD estimate can also be high, since it is quadratic in the number of generated sequences. A practical workaround is to reduce the number of samples or temperature values used when computing the MMD trajectory. However, the stability and reliability of MMD-Flagger under such reduced settings remain an open question.
Broader Impact. Our work contributes to the development of safer and more widespread applications of LLMs. Indeed, MMD-Flagger can enhance translation-related tasks by identifying hallucinations in LLMs outputs.
5 Conclusion
In this work, we introduced MMD-Flagger, a method for hallucination detection based on the Maximum Mean Discrepancy (MMD) metric. The key idea is to compare output sequences generated by beam search and stochastic sampling across varying temperature settings. By measuring the distributional distance between these sequences using MMD, MMD-Flagger achieves detection performance on par with the strongest uncertainty-based estimation methods on two machine translation benchmarks. The method’s performance is robust with respect to changes in the output vectorization and choice of the kernel.
Future work includes extending MMD-Flagger to other generation tasks such as summarization or question-answering, using LLMs such as Qwen (Yang-2024-Qwen, ) or Llama (Grattafiori-2024llama, ) as models to inspect. Benchmark datasets are available (Maynez-etal-2020-faithfulness, ; Bao-2025-naacl, ), and the sampling strategy is the same. The challenge here is the size of the models at hand and the outputs.
Another promising research direction is to investigate the interpretability of MMD-Flagger. One possibility is to leverage recent work in feature importance for MMD-based testing sutherland2017 ; mitsuzawa-2023 . This would allow to pinpoint embedding features responsible for the detection of hallucinated content by MMD-Flagger. Once such variables are located in intermediate layers, interpretability techniques such as Layerwise Relevance Propagation (Montavon-2019, ) could be applied to analyze how these variables relate to specific encoder layers or segments of the input.
Acknowledgements
This work has been supported by the French government, through the NIM-ML project (ANR-21-CE23-0005-01). We extend our sincere thanks to Prof. Motonobu Kanagawa for insightful discussions and valuable feedbacks to the conceptualization of this work. We thank Dr. Jean-Flavien Bussotti and Ms. Shilpa Sunil Vishwakarma for their careful proofreading and helpful suggestions to improve the clarity and readability.
Appendix
This appendix is organized in the following way: Appendix A describes relabeling procedure of annotation labels of LFAN-HALL dataset. Appendix B details the experimental setup. Appendix C contains additional results related to the hallucination detection experiments (Section 3 of the main paper). Key facts regarding the datasets used for experiments are collected in Section D. Appendix E shows qualitative results of the hallucination detection by MMD-Flagger. Comparing kernel length settings are described in Appendix F. We show the MMD-Flagger hallucination detection results using vectors from intermediate layers of models in Appendix G. Finally, we investigate the variability of the proposed hallucination detection procedure in Appendix H.
[…]
Contents of the Appendix
Appendix A LFAN-HALL Dataset Modification and Label Correction
Some annotation labels in the LFAN-HALL dataset appear to be inconsistent with the definitions given by the paper. Referring to the annotation guideline provided [27] [^1] https://github.com/deep-spin/hallucinations-in-nmt/tree/main/data, the strongly detached label is applied when “there is some content that is supported by the source sentence. However, there is a significant part of the translation that is not (it conveys different meaning to the meaning in the source sentence). Importantly, the content that is hallucinatory does not reflect a genuine effort to translate any part of the source sentence”. In Table 3, we report an example which, in the original datasets, has two labels: omission (of the machine translation error) and strongly detached (of the hallucination). This example is considered to be labeled as the omission (of the machine translation) since the machine translation correctly translated the phrase of the source text A proud celebration, yet the machine translation lacks the phrase that gives cause for celebrating. Therefore, this example is not a case of hallucination.
Table 3: An example of modified annotation labels in LFAN-HALL dataset. The original annotation labels are omission and strongly detached. See the description above.
Source text (German) Reference translation Machine translation
Ein stolzes Jubiläum, das zum Feiern Anlaß gibt. A proud anniversary that gives cause for celebrating. A proud celebration.
We reviewed the annotation labels, focusing on the translations originally labeled as hallucinations, that is, those marked as “strongly detached”, “fully detached”, or “repetition.” Table 4 presents a confusion matrix comparing our verified labels against the original annotation. Notably, substantial discrepancies were found in cases labelled as a combination of “omission” and “strongly detached.” Specifically, 65% (19 examples) of those were re-labelled as “omission,” which is not categorised as hallucination. This may suggest that the annotators have misinterpreted the definition of “strongly detached.” In particular, when a machine translation contains an omission error, the resulting output may appear distant from the source text, potentially leading to misclassification.
Table 4: Confusion matrix comparing verified and original annotation labels in the LFAN-HALL dataset. Cell values indicate the frequency of each annotation label. The symbol “&” denotes the co-occurrence of two error types. Abbreviations are as follows: “NE”: named entity translation error; “omit”: omission error; “rep”: repetition-based hallucination; “full”: fully detached hallucination; “strong”: strongly detached hallucination. Columns are not shown when the frequency is zero. Yellow-highlighted cells denote notable discrepancies between the verified and original annotation labels.
[…]
full 117 0 0 0 0 0 0 0 4
full & /NE 0 1 0 0 0 0 0 0 0
full & rep 0 0 11 0 0 0 0 0 0
NE 0 0 0 2 0 0 0 0 1
NE & omit 0 0 0 0 0 0 0 0 1
NE & strong 0 0 0 0 0 0 0 0 0
omit 0 0 0 1 0 19 0 0 1
omit & rep 0 0 0 0 1 0 0 0 0
omit & strong 0 0 0 0 0 1 0 0 0
rep 0 0 0 0 0 0 29 0 0
rep & strong 0 0 0 0 0 0 0 45 2
strong 0 0 0 1 0 8 0 0 69
none 0 0 0 0 0 1 1 0 8
Appendix B Experimental Setup
B.1 Computational Environment and Implementation Details
Experiments are done on the following computational environment: Xeon Platinum 8268 Processor with a single GPU of Tesla V100S-PCIE $32$GB and $755$GB RAM.
B.2 Translation models
Model for LFAN-HALL. The translation model is provided by the authors [27] [^2] https://github.com/deep-spin/hallucinations-in-nmt. The model is trained by the authors using the fairseq package on the WMT18 DE-EN dataset [8]. The model is a Transformer model consisting of hidden vectors of $512$ dimensions, $6$ layers for both the encoder and decoder with $8$ attention heads. The total parameter size of the model is approximately $77$M parameters.
The translation speed is roughly $2$ seconds to sample $25$ sequences. In the experiments using vectors from the intermediate layers in Appendix G, the translation speed is significantly slower: approximately $30$ seconds for sampling $25$ sequences. This is because of the necessity of a custom codebase for extracting the vectors from the intermediate layers. Since we, regardless of our best efforts, could not find interfaces of the fairseq package to access and extract the intermediate layers during the inference mode, it is necessary to implement custom code, which causes a significant slowdown.
Model for Halomi. The translation model is the multilingual translation model [10], and the model is available on the HuggingFace Hub [^3] https://huggingface.co/facebook/nllb-200-distilled-600M. The model is based on the multilingual encoder-decoder model designed for the machine translation task [^4] The base model in the transformers package is M2M100: https://huggingface.co/docs/transformers/en/model_doc/m2m_100. The model consists of $12$ layers, and each layer generates a vector of $1,024$ dimensions.
In this case, it takes approximately $4$ seconds to sample $25$ sequences.
B.3 Implementation details
MMD-Flagger. We use an open source codebase to compute the MMD distances [37] [^5] https://github.com/Kensuke-Mitsuzawa/mmd-two-sample-test-variable-selection. The codebase is implemented using PyTorch [42]. About the computational speed, MMD-Flagger requires around $1$ second for computing MMD distances of $10$ temperature parameters when the vector aggregation mode is concat and the dimension size is $D=53,760$ and the sample size of $H_{\text{sto}}^{\tau}$ is $25$. Therefore, the required time for computing an MMD distance is around $0.1$ seconds per $\tau$ parameter.
MC-Dsim. We activate the dropout of the model for simulating the stochastic process and obtaining the translation token sequences. For calculating a distance between the token sequences using the drop-out and without, we use the METEOR metric [3] of the nltk package [7] to compute the similarity score [^6] https://www.nltk.org/api/nltk.translate.meteor_score.html.
Seq-LogProb. fairseq and transformers packages both give the log probability for a generated sequence.
Appendix C Additional experimental results
C.1 Additional results for the Halomi Dataset
Table 5 represents the numbers of true-positive and false-positive per annotation label.
Table 5: Detection performance per error label on the Halomi dataset. The first row represents short names of hallucination labels that represent the detachment levels of the translation text, and the second row is the total count of the annotation label. The first column shows detection methods description of which is in Section 3.3. The value formats are “true positive / false positive.” We highlight the highest values of true positive in each annotation label.
small partial full
Count 225 291 218
MMD-Flagger (avg) 187 / 1,938 239 / 1,886 180 / 1,945
MMD-Flagger (concat) 187 / 1,989 244 / 1,932 174 / 2,002
SeqLogProb 65 / 958 142 / 881 73 / 950
MC-Dsim 109 / 836 158 / 787 139 / 806
TNG 2 / 0 2 / 0 2 / 0
Table 6 shows the detection performance for each source and target language pair. Among the $18$ language pairs analyzed, we can categorize them into three types based on the availability of training resources for the translation model: low-resource, high-resource, and zero-shot languages. Among the nine languages studied, the following three are considered low-resource languages when training the translation model: kas-Deva, mni-Beng, and yor-Latn. The source and target language pairs between Spanish and Yoruba are classified as a zero-shot learning setting that generates the translation without the training data of the corresponding source and target pair. The remaining pairs consist of high-resource languages. For detailed training information, please refer to [11].
We posited that the detection performance by MMD-Flagger may correlate with the volume of training data: low, high, and zero-shot. However, as evidenced by the results presented in Table 6, the performance of MMD-Flagger in hallucination detection does not exhibit a significant relationship with the amount of data on which the model was trained. Notably, when comparing the recall scores for MMD-Flagger, the high-resource language pairs, specifically “spa-Latn to eng-Latn” yields a lower recall score. Conversely, the recall scores for zero-shot language pairs, such as yor-Latn and spa-Latn, demonstrate relatively more favorable results.
Table 6: Detection performance per language pair on the Halomi dataset. The recall and precision scores are shown in the format of “recall (precision)”. Language pairs are written with language codes of the dataset [11]. The sub-section names “High”, “Low”, “Zero-shot” represent the training data volume for training the translation model [11]. We highlight the best recall score per language pair.
Source and Target Langs Count MMD-F(avg) MMD-F(concat) SeqLogProb TNG
High-Resource Languages
arb-Arab to eng-Latn 156 0.67 (0.17) 0.75 (0.18) 0.42 (0.16) 0.00 (0.00)
deu-Latn to eng-Latn 155 0.93 (0.13) 0.87 (0.12) 0.53 (0.13) 0.00 (0.00)
eng-Latn to arb-Arab 144 0.88 (0.06) 0.75 (0.05) 0.12 (0.02) 0.00 (0.00)
eng-Latn to deu-Latn 146 0.86 (0.12) 0.79 (0.11) 0.43 (0.10) 0.00 (0.00)
eng-Latn to rus-Cyrl 146 1.00 (0.05) 0.80 (0.04) 0.40 (0.03) 0.00 (0.00)
eng-Latn to spa-Latn 153 0.73 (0.22) 0.68 (0.21) 0.00 (0.00) 0.00 (0.00)
eng-Latn to zho-Hans 160 0.69 (0.18) 0.59 (0.14) 0.55 (0.25) 0.03 (1.00)
rus-Cyrl to eng-Latn 158 0.92 (0.09) 0.83 (0.09) 0.17 (0.03) 0.00 (0.00)
spa-Latn to eng-Latn 160 0.58 (0.22) 0.55 (0.20) 0.52 (0.27) 0.00 (0.00)
zho-Hans to eng-Latn 159 0.84 (0.21) 0.81 (0.19) 0.44 (0.22) 0.03 (1.00)
Low-Resource Languages
eng-Latn to kas-Deva 184 0.90 (0.42) 0.89 (0.41) 0.38 (0.38) 0.00 (0.00)
eng-Latn to mni-Beng 197 0.82 (0.60) 0.90 (0.63) 0.49 (0.73) 0.00 (0.00)
eng-Latn to yor-Latn 195 0.90 (0.16) 0.90 (0.16) 0.41 (0.15) 0.00 (0.00)
kas-Deva to eng-Latn 151 0.79 (0.40) 0.77 (0.39) 0.47 (0.48) 0.00 (0.00)
mni-Beng to eng-Latn 152 0.89 (0.80) 0.85 (0.79) 0.45 (0.80) 0.00 (0.00)
yor-Latn to eng-Latn 146 0.82 (0.16) 0.91 (0.16) 0.45 (0.17) 0.00 (0.00)
Zero-shot Languages
yor-Latn to spa-Latn 152 0.85 (0.51) 0.85 (0.47) 0.00 (0.00) 0.00 (0.00)
spa-Latn to yor-Latn 151 0.83 (0.35) 0.87 (0.35) 0.33 (0.30) 0.00 (0.00)
C.2 Additional results for the LFAN-HALL Dataset
More intuition on MMD-Flagger. Figure 3 illustrates examples of the intuition of detecting the hallucination by MMD-Flagger.
In detecting the hallucination by MMD-Flagger, the MMD distance at $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is the principal factor of forming the U-shape in the MMD distances trajectory, as shown in Figure 1; the U-shape is formed since $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is larger than $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ values. Figure 3 depicts distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ per ground-truth annotation label. These distributions represent clear differences by the annotation labels; Comparing these distributions, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ with the hallucination label represents the high volume in the center, while the distribution with the correct label depicts the high volume at the quite low $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ value. In case of the machine translation errors, the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ looks quite different from the distribution of the correct label. There may be room for employing MMD-Flagger for detecting the machine translations; However, further studies are required.
Figure 3: Distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ by annotation label. The horizontal axes represent values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$, and the vertical axes are percentages. Comparing two distributions “Correct” and “Hallucination”, we observe frequent larger values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ in “Hallucination”. This tendency suggests that the U-shape of the trajectory is likely to form.
Figure 4 displays the distributions of the $\tau$ values at which the trajectory $\mathcal{T}_{\widehat{MMD}}$ reaches its minimum. In the case of the hallucination label, the minimum MMD values are often observed at $\tau>0.2$, indicating that the $\mathcal{T}_{\widehat{MMD}}$ trajectory tends to exhibit a U-shape, as exemplified in Figure 1.
Figure 4: Distributions of $\arg\min(\mathcal{T}_{\widehat{MMD}})$. The horizontal axes represent $\tau$, and the vertical axes are the percentages. Comparing two distributions “Correct” and “Hallucination”, the high frequency of being $\arg\min(\mathcal{T}_{\widehat{MMD}})$ is at $0.1$ in the “Correct” label, while $>0.1$ in the “Hallucination” label. This pattern suggests that the U-shape is likely to form in the “Hallucination” label.
Detection Performance per Label. Table 7 represents the hallucination detection performance per annotation label. The kernel settings of the MMD-Flagger are the same as described in Section 3.2.
Table 7: Detection performance per error label on the LFAN-HALL dataset. The first row represents short names of hallucination labels that represent the detachment levels of the translation text, and the second row is the total count of the annotation label. The hallucination labels are full-d: “full-detachment”, strong-d: “strong-detachment”, rep: “repetition”. The labels of omission and NE (Named Entity) are machine translation errors. The value formats are “true positive / false positive”. We highlight the highest true positive values for each annotation label.
full-d strong-d rep omission NE
Counts 133 131 88 198 41
MMD-Flagger (avg) 115 / 1,739 102 / 1,752 69 / 1,785 117 / 1,737 32 / 1,822
MMD-Flagger (concat) 118 / 1,748 108 / 1,758 69 / 1,797 125 / 1,741 31 / 1,835
MC-DSim 107 / 1,178 110 / 1,175 66 / 1,219 85 / 1,200 18 / 1,267
Seq-Logprob 122 / 1,163 102 / 1,183 63 / 1,222 74 / 1,211 22 / 1,263
TNG 1 / 19 17 / 3 20 / 0 0 / 20 0 / 20
Threshold Selection. Figure 5 represents the uncertainty score distributions and Receiver operating characteristic (ROC) curves of Seq-Logprob and MC-DSim.
(a) Seq-Logprob
(b) MC-DSim
Figure 5: (Top) Score distributions of Seq-Logprob (left) and MC-DSim (right) on LFAN-HALL dataset. The horizontal axis represents an evaluation score for a datapoint, and the vertical axis represents the frequency. The vertical red dashed line is the threshold for the score, selected at $40$-th percentile of the score distribution. (Bottom) ROC curves of Seq-Logprob (left) and MC-DSim (right) on LFAN-HALL dataset. “AUC” in the caption stands for the Area Under the Curve. The orange dashed lines indicate the chance level.
Appendix D Datasets
The dataset description is found at Section 3.1. In this section, we describe additional information about the datasets and label statistics.
D.1 LFAN-HALL Dataset [27]
The original dataset file is provided by the authors [^8] https://github.com/deep-spin/hallucinations-in-nmt. As we discuss in Appendix A, we reviewed the annotation labels and modified it. In the following section, we show label statistics of the modified dataset.
[…]
Table 8: Label distribution of the LFAN-HALL dataset. Labels are classified into three categories: no-error, hallucination, and mt-error. The comma “,” represents that the two labels are given.
Label Count Ratio
no-error 2,885 0.8455
hallucination 291 0.0852
mt-error 233 0.0682
hallucination,mt-error 3 0.0008
Table 9: Label distribution of the LFAN-HALL dataset. The comma “,” represents the two labels are given.
Label Count Ratio
no-error 2,885 0.8455
error-omission 193 0.0565
error-full 121 0.0354
error-strong 83 0.0243
error-repetitions, error-strong 47 0.0137
error-named-entities 37 0.0108
error-repetitions 29 0.0084
error-full, error-repetitions 11 0.0032
error-named-entities, error-omission 3 0.0008
error-omission, error-strong 1 0.0002
error-omission, error-repetitions 1 0.0002
error full, error-named-entities 1 0.0002
D.2 Halomi Dataset [11]
The label statistics of the Halomi dataset are shown in Table 10.
Table 10: Hallucination label distribution of the Halomi dataset.
Class label Count Ratio
1 No hallucination 2,131 0.74
3 Partial hallucination 291 0.10
2 Small hallucination 225 0.07
4 Full hallucination 218 0.07
Appendix E Qualitative results
In this section, we show examples of detections; the source text and translations are in the left tables, and the corresponding MMD trajectories $\mathcal{T}_{\widehat{MMD}}$ are in the right plots. The examples are all from the LFAN-HALL dataset, and the source language is German. In each table, we show one example of the stochastic sampled translation $y_{\text{sto}}^{\tau=0.1}$.
The detection examples are all with the MMD-Flagger with the vector representation of avg and the kernel setting of gaussian using the $25$-th percentile value for computing the Gaussian kernel length scale, reported in Table 2.
E.1 Correct hallucination detections (True Positive)
Source Nur der Stuck an der Decke uns das Wandabendmahl konnten nicht restauriert werden.
Reference Only the stuccos of the ceiling and the wall could not be restored.
$y_{\text{beam}}$ The room was very small and the bathroom was very small.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in a quiet area of the city.
Source Auf den jeweiligen Dorfplätzen finden sich Alt und Jung zum Schwätzchen und zum Feiern zusammen.
Reference On both places, young and old meet each other to have a chat or to celebrate together.
$y_{\text{beam}}$ Old and young people will find themselves together in the village’s respective squares for fun and fun.
$y_{\text{sto}}^{\tau=0.1}$ The village’s squares are home to old and young people who want to swim and celebrate.
Source Vermutlich überall ist Situation mit Taxi so - wenn man einen Fremdsprachepreis für ein Taxi geht oben spricht.
Reference Probably everywhere situation with taxi is like this - if one speaks a foreign language price for a taxi goes up.
$y_{\text{beam}}$ The hotel is located in the heart of the city, just a few steps away from the train station.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in the heart of the city, just a few steps from the beach and the beach.
E.2 Correct translations (True Negative)
Source Zum Abschluss der Aussprache stellte der Präsident Folgendes fest:
Reference The president concluded the debate as follows:
$y_{\text{beam}}$ At the end of the debate, the President said:
$y_{\text{sto}}^{\tau=0.1}$ At the end of the debate, the President noted that:
Source Politiken, auf die besonderes Augenmerk zu richten ist, und
Reference policies which should be given special attention, and
$y_{\text{beam}}$ policies to which particular attention needs to be paid; and
$y_{\text{sto}}^{\tau=0.1}$ policies to which particular attention must be paid; and
E.3 Detection by error (False Positive)
False Positive cases are often observed when $y_{\text{beam}}$ and $y_{\text{sto}}^{\tau=0.1}$ are semantically similar; however, dissimilarity due to different expressions or words. These false-positive errors would be due to the vector representations avg of the word embedding vectors. Due to the aggregation, the Gaussian kernel failed to compute the similarity correctly, and thus, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ becomes a high value.
Source Ich komme zum Schluß noch darauf zu sprechen.
Reference I shall come back to this point again at the end.
$y_{\text{beam}}$ Finally, I would like to mention this.
$y_{\text{sto}}^{\tau=0.1}$ I will end by mentioning this.
Source Wollen sie eine Firma gründen oder umstrukturieren?
Reference Do you intend to market new product or service?
$y_{\text{beam}}$ Do they want to set up or restructure a company?
$y_{\text{sto}}^{\tau=0.1}$ Do you want to start a company or restructure it?
E.4 Fail to detect (False Negative)
The false-negative errors are often observed when the stochastic sampled translations $y_{\text{sto}}^{\tau=0.1}$ are really similar to the beam search translation $y_{\text{beam}}$. In this case, $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ becomes a small value and is hard to form a U-shape; therefore, MMD-Flagger failed to detect the hallucination.
Source Februar 1997 angenommenen Entschließung ABl.
Reference 61. adopted on 19 February 1997;
$y_{\text{beam}}$ Resolution OJ L 208, p.
$y_{\text{sto}}^{\tau=0.1}$ Resolution OJ L 208, p.
Source Oktober zu verlängern.
Reference police and penitentiary.
$y_{\text{beam}}$ The Commission has decided to extend the programme.
$y_{\text{sto}}^{\tau=0.1}$ The Commission has decided to extend the programme.
Source Herr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit
Reference Martin BARTENSTEINFederal Minister for Economic Affairs and Labour
$y_{\text{beam}}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit.
$y_{\text{sto}}^{\tau=0.1}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit
Appendix F Comparison of Kernel Length Scale Selection
The “median heuristic” [23] is a widely used method in the kernel literature for selecting the length scale of a kernel function. It computes the pairwise distance matrix (e.g., using the Euclidean distance) over a set of data points and uses the median of these distances as the length scale. While the $50$-th percentile (i.e., the median) is commonly used, other percentiles can also be considered.
In Table 11, we evaluate hallucination detection performance under different configurations of percentile values and length scale shapes (single or vector).
Table 11: Comparison of hallucination detection performance by configuration of a Gaussian kernel function. The vector representation is by the word embedding layer, and vector aggregation is by concatenating vectors of $T$ tokens.
Aggregation Length scale Percentile LFAN-HALL Halomi
concat single 12.5 0.840 (0.132) 0.824 (0.278)
25 0.850 (0.134) 0.824 (0.278)
50 0.837 (0.133) 0.823 (0.279)
62.5 0.830 (0.133) 0.824 (0.280)
75 0.837 (0.134) 0.820 (0.279)
87.5 0.840 (0.136) 0.820 (0.279)
avg single 12.5 0.816 (0.129) 0.826 (0.285)
25 0.816 (0.129) 0.825 (0.281)
50 0.813 (0.130) 0.820 (0.285)
62.5 0.813 (0.130) 0.817 (0.284)
75 0.810 (0.129) 0.814 (0.284)
87.5 0.810 (0.129) 0.809 (0.283)
Appendix G Hallucination Detection using Vectors from Intermediate Layers
In this section, we report MMD-Flagger detection scores using vectors at intermediate layers of a model. Previous work reports that the model’s intermediate hidden states correlate with the existence of hallucinated token sequences [11, 2, 49]. Thus, vector representations from the intermediate vector values of the outputs could also contain valuable information for the hallucination detection task.
In Section 2.2, we introduce the vector representation using the word embedding matrix $E\in\mathbb{R}^{V\times D}$. Instead of employing the matrix $E$, we access an intermediate layer of the LLM’s decoder and extract vector representations after the “Add & Norm” operation at each layer, as defined in the Transformer architecture [53], where “Add & Norm” means the residual connection followed by layer normalization. In this work, the model for the LFAN-HALL dataset has $6$ layers (the dimension size of $512$), and the model for the Halomi dataset has $12$ layers (the dimension size of $1,024$).
We use the Gaussian kernel for the MMD-Flagger. For calculating the length scale, we collect the vector representations of $200$ calibration data and select the $25$-th percentile of $L^{2}$-distance, in the same manner described in Section 3.2.
Table 12 shows the recall and precision scores. The recall scores’ tendencies of two datasets indicate differently. On the LFAN-HALL dataset, the best recall score is at the $1$-th layer, and the recall score gradually becomes lower as the layer goes deeper. On the Halomi dataset, on the contrary, the recall score is low at the earlier layers, nevertheless, the recall score becomes higher as the layer goes deeper. These tendencies may explain a specific layer of the LLM causing the hallucinations; for instance, the first layer of the LFAN-HALL model lets MMD-Flagger plot the U-shape trajectories more often; therefore, the vector representation of the first layer may explain the cause.
Table 12: Comparison of hallucination detection performance by configuration of a Gaussian kernel function. The vector representation is by the word embedding layer, and vector aggregation is by concatenating vectors of $T$ tokens. The symbols “–” represents the absence of corresponding layers in the model.
Layer Number Aggregation Kernel LFAN-HALL Halomi
layer 1 avg Gaussian 0.891 (0.104) 0.676 (0.310)
layer 2 avg Gaussian 0.884 (0.103) 0.673 (0.310)
layer 3 avg Gaussian 0.874 (0.103) 0.681 (0.305)
layer 4 avg Gaussian 0.857 (0.101) 0.685 (0.303)
layer 5 avg Gaussian 0.850 (0.103) 0.694 (0.304)
layer 6 avg Gaussian 0.840 (0.103) 0.697 (0.303)
layer 7 avg Gaussian 0.720 (0.301)
layer 8 avg Gaussian 0.727 (0.296)
layer 9 avg Gaussian 0.736 (0.348)
layer 12 avg Gaussian 0.767 (0.264)
layer 1 concat Gaussian 0.918 (0.104) 0.702 (0.285)
layer 2 concat Gaussian 0.907 (0.101) 0.695 (0.287)
layer 3 concat Gaussian 0.901 (0.100) 0.691 (0.285)
layer 4 concat Gaussian 0.897 (0.100) 0.699 (0.283)
layer 5 concat Gaussian 0.890 (0.099) 0.717 (0.282)
layer 6 concat Gaussian 0.894 (0.101) 0.747 (0.283)
layer 7 concat Gaussian 0.783 (0.282)
layer 8 concat Gaussian 0.816 (0.270)
layer 9 concat Gaussian 0.853 (0.266)
layer 10 concat Gaussian 0.875 (0.262)
layer 11 concat Gaussian 0.895 (0.257)
layer 12 concat Gaussian 0.914 (0.263)
Appendix H Detection Stability of MMD-Hallucination-Flagger
In this section, we discuss the stability of MMD-Flagger in detecting hallucinations, focusing on the impact of randomness and sample size. The detection process involves stochastic sampling, where a random selection of the next token is made based on the probability distribution generated by the softmax function, as detailed in Eq.(1). Our empirical findings suggest that MMD-Flagger is effective in identifying hallucinations with sequences comprising just $25$ tokens.
We used a subset of the LFAN-HALL dataset to iteratively generate token sequences, performing this process 10 times with sample sizes of $\{10,25,50,100\}$. This subset consists of translation examples that all have hallucination labels, totaling $294$ examples; consequently, we only report the recall score.
The vector representation is set to avg, and the kernel employed is a Gaussian kernel. The length scale is determined to be the $25$-th percentile of $200$ calibration data, following the same method described in Section 3.2. The settings for the Gaussian kernel remain consistent across all experiments, ensuring that the computed MMD distances are comparable.
Figure 6 illustrates the bar plots representing recall scores. Notably, the range of recall scores for a sample size of $25$ is closely aligned with that of $100$ samples. This observation suggests that a sample size of $25$ is adequate for effectively detecting hallucinations.
Figure 6: Detection performance of MMD-Flagger using various sample sizes. The horizontal and vertical axes represent the number of samples and the recall score. The bars represent the average of recall scores in $10$ repetitions, and the error bars are the variance.
Figure 7 illustrates the MMD trajectory, denoted as $\mathcal{T}_{\widehat{MMD}}$, based on $10$ repetitions with varying sample sizes of $\{10,25,50,100\}$.
The trajectory $\mathcal{T}_{\widehat{MMD}}$ consistently exhibits the U-shape, regardless of the sample size. Particularly, the variance, represented by the error bars, decreases with increasing sample size. A trajectory derived from a sample size of $100$ is expected to closely align with the convergence line, indicating a more stable representation of the MMD trajectory.
The MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ plays a significant role in shaping the U-curve of $\mathcal{T}_{\widehat{MMD}}$. Therefore, we analyze the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ across 294 examples.
Our assumption is that if $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ is close to the converged value, the distance value does not change significantly, regardless of the random seeds used or the sample size. As a result, the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ are expected to exhibit similar shapes.
Figure 8 illustrates the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ at the first and third repetitions. With a sample size of 10, the two distributions appear quite different; however, at a sample size of 100, the distribution shapes seem to converge.
(a) With $10$ samples
(b) With $25$ samples
(c) With $50$ samples
(d) With $100$ samples
Figure 7: Comparing $\mathcal{T}_{\widehat{MMD}}$ of sample sizes $\{10,25,50,100\}$. The line is the average value of $10$ repetitions, and the area plots are the variance values.
(a) At 1st repetition using $10$ samples
(b) At 3rd repetition using $10$ samples
(c) At 1st repetition using $25$ samples
(d) At 3rd repetition using $25$ samples
(e) At 1st repetition using $50$ samples
(f) At 3rd repetition using $50$ samples
(g) At 1st repetition using $100$ samples
(h) At 3rd repetition using $100$ samples
Figure 8: Comparisons of distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto})$ over $294$ examples. The left column shows the distributions at first repetition, and the right column is for the one of the third repetition.
Appendix I MMD trajectory filtering
This section introduces the hallucination detection performance after applying the trajectory filtering to $\mathcal{T}_{\widehat{MMD}}$. However, we do not observe substantial differences by applying trajectory filters.
The MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ can exhibit irregular shapes, as illustrated in Figure 9 and filtered trajectories by the average window. To address these irregularities, a smoothing method can be applied to create a more uniform trajectory. By utilizing smoothing techniques, the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ becomes easier to manage, which may enhance the ability of MMD-Flagger to detect hallucinations more effectively.
Figure 9: An example of the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ and filtered trajectories with the average window by $2$ and $3$ sizes. The blue solid line is the original MMD trajectory. The purple dashed line is the filtered line with the window size of $2$, and the green dashed line is the filtered line with the window size of $3$. The red circle is the minimum value of the original MMD trajectory, and black “x” marks are the minimum values on the filtered trajectory.
Table 13 compares the hallucination detection performance of the original MMD trajectory and filtered trajectories. The length scale computation at $25$-th percentile; same as Section 3.2.
Table 13: Comparing the hallucination detection of the original MMD trajectory and filtered trajectories. The score formats are “recall (precision)”.
Vector representation Kernel Filtering Window size LFAN-HALL
avg dot No Filter 0.820 (0.125)
dot Average mean 2 0.782 (0.143)
dot Average mean 3 0.734 (0.145)
gaussian No Filter 0.823 (0.129)
gaussian Average mean 2 0.785 (0.131)
gaussian Average mean 3 0.737 (0.133)