### v3

# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

Abstract
Large language models (LLMs) have become pervasive in our everyday life. Yet, a fundamental obstacle prevents their use in many critical applications: their propensity to generate fluent, human-quality content that is not grounded in reality. The detection of such hallucinations is thus of the highest importance. In this work, we propose a new method to flag hallucinated content: MMD-Flagger. It relies on Maximum Mean Discrepancy (MMD), a non-parametric distance between distributions. On a high-level perspective, MMD-Flagger tracks the MMD between the output to inspect and counterparts generated with various temperature parameters. We show empirically that inspecting the shape of this trajectory is sufficient to detect most hallucinations. This novel method is benchmarked on machine translation and summarization datasets, on which it exhibits competitive performance relative to natural competitors.
MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations
Kensuke Mitsuzawathanks: Corresponding author. Université Côte d’Azur, CNRS, LJAD, France kensuke.mitsuzawa@unice.fr          Damien Garreau Center for Artificial Intelligence and Data Science (CAIDAS) Julius-Maximilans-Universität Würzburg, Germany damien.garreau@uni-wuerzburg.de
1 Introduction
Neural probabilistic modelling in an auto-regressive framework (Bengio et al., 2003), powered by Transformers (Vaswani et al., 2017), underpins modern NLP, enabling advances in translation (Kocmi et al., 2024), summarization (Shakil et al., 2024), question answering (Daull et al., 2023), and more. Large Language Models (LLMs) such as GPT (Radford et al., 2018), popularized by ChatGPT and now used by millions (Deng et al., 2023), can address diverse tasks such as the machine translation or summarization. But it can suffer from hallucinations, fluent yet factually incorrect generated content (Ji et al., 2023). Though relatively infrequent (Guerreiro et al., 2023b), hallucinations are especially concerning in translation and summarization, where outputs may deviate semantically from the source (see Table 1), and in high-stakes domains such as law (Benkirane et al., 2024), where even a single error can erode trust. Given that detecting hallucinations at the test time remains a major challenge (Sriramanan et al., 2024), we focus on detection in white-box settings and inspecting a single LLM’s output. This allows us to retain full internal access to the model’s parameters and intermediate states.
The emergence of hallucinations in generated output is involved with the decoding strategy used for next-token prediction (Shi et al., 2024). While no universally optimal strategy exists, beam search (Meister et al., 2020) and its variants are often practical choices, particularly for models fine-tuned through alignment with human feedback (Shi et al., 2024). However, beam search and its variants are prone to the so-called likelihood trap (Das and Balke, 2022), where sequences with high probability can nevertheless yield poor-quality or uninformative outputs, which can cause the hallucination.
An alternative decoding strategy is to sample the next token directly from the probability distribution given by the model’s softmax layer. The key advantage of this approach is its stochasticity, which allows diverse outputs to be generated from the same input. A common refinement introduces a temperature parameter ($\tau$), which adjusts the sharpness of the probability distribution. Stochastic sampling, therefore, provides opportunities to improve generation quality under appropriate parameter choices (Shi et al., 2024).
Table 1: Examples of hallucinated translations from the LFAN-HALL dataset generated by beam search ($y_{\text{hyp}}$), and stochastic samples with varying temperature parameters ($y_{\text{sto}}^{\tau}$). The MMD distances between $y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.5}$ as a function of $\tau$ can be visualized as the “Hallucinated” in Figure 1.
Source (German) Gutes Angebot bei Techno, House und Minimal Techno.
Reference translation Good choice of Techno, House and Minimal Techno tracks.
$y_{\rm hyp}$ The staff were very friendly and helpful. The room was clean and comfortable.
$y^{\tau=0.1}_{\rm sto}$ The hotel is very close to the city centre and the train station.
The hotel is very well located, close to the beach and the beach.
$y^{\tau=0.5}_{\rm sto}$ The staff was very helpful and friendly.
The hotel is very close to the city centre and the metro station.
$y^{\tau=0.8}_{\rm sto}$ It had the best overall prices. It was close to the beach and business.
The breakfast buffet was very good häufude and the location is ideal for getting to the town.
The central idea of this study is to inspect a hypothesis output, generated with a default decoding strategy, with stochastic samples obtained under varying temperature. At low temperature settings, we observe that the sampled outputs are highly similar to the hypothesis if the model is not hallucinating, and moderately dissimilar if it is. We believe this occurs because of the following reason: when a clear optimal candidate token exists, the default decoding strategy and stochastic sampling tend to favor the most probable token. However, in cases where no single candidate clearly dominates (often indicative of hallucination), the outputs are more likely to diverge. In contrast, at high temperature, generation becomes increasingly random, and sampled outputs tend to be substantially different from the default output regardless of the model’s behavior. This behavior is illustrated in Table 1. As a consequence, plotting the similarity between the default output and the stochastic outputs yields two distinct curves: either a monotonously increasing trajectory in the non-hallucinating scenario, or a $U$-shape curve in the hallucinating scenario. This can be visualized in Figure 1.
Figure 1: MMD-Flagger at work on the LFAN-HALL dataset: two outputs, generating two MMD trajectories. In orange, the trajectory for the hallucinated translation (see Table 1), in maroon, the trajectory corresponding to the correct one. Text boxes show examples of $y_{\rm hyp}$ and samples from $Y_{\rm sto}^{\tau}$ (two of 25 samples), generated by stochastic decoding with temperatures $\tau$. We flag $y_{\rm hyp}$ as a hallucination when the minimum value of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ (the vertical black dashed line) falls within the blue region. Otherwise, the output is deemed plausible (minimum falling into the blue region).
Drawing on these findings, we introduce a novel hallucination detection method: MMD-Flagger. The key idea of the method is to simply i) plot the similarity between default output and stochastic samples as a function of the temperature, and ii) automatically detect whether this curve is monotonously increasing or $U$-shaped. In the latter case, we flag the output as hallucination. As a measure of similarity, we propose to use the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012). As a shape detection mechanism, we simply estimate the inflection point of the trajectory and check whether it is located near 0 (see Figure 1).
The paper is organized as follows. In Section 2, we detail the operating procedure of MMD-Flagger after recalling the necessary background on temperature sampling and MMD. We benchmark our method against several natural competitors in Section 3, before concluding in Section 5. The code for all experiments is provided in code.tar.gz and will be publicly available after publication.
1.1 Related work
Hallucination detection in LLMs has attracted significant attention, with a variety of methods proposed in recent years. We follow the categorization outlined by Park et al. (2025) and highlight the approaches most relevant to our work.
Consistency-based methods. These methods detect hallucinations by comparing multiple generations from the same prompt under different decoding strategies. Our proposed method, MMD-Flagger, belongs to this category. MC-DSIM (Fomicheva et al., 2020; Zerva et al., 2021) leverages Monte-Carlo dropout (Gal and Ghahramani, 2016) to generate multiple outputs and then computes an uncertainty score using a similarity metric in natural language space. Semantic Uncertainty Nikitin et al. (2024) similarly collects stochastically sampled outcomes, builds a graph representation, and derives an uncertainty score from graph structure. SelfCheckGPT (Manakul et al., 2023) compares candidate generations against stochastically sampled alternatives. SelfCheckGPT and Semantic Uncertainty require external models (e.g. inference models) for building a graph or scoring. In contrast, MMD-Flagger operates without external resources.
Logits-based methods. These approaches use log-probabilities to assess the fluency of the generated token sequence (Ren et al., 2023). However, hallucinations that are syntactically plausible can still receive high log-probabilities, making them difficult to detect. In addition, setting thresholds for logits-based scores presents practical challenges.
Hidden-State-based methods. Several works exploit internal representations to identify hallucinations. Cross-attention patterns in encoder–decoder models have been used for detection (Guerreiro et al., 2023a; Ferrando et al., 2022; Dale et al., 2023), though such methods are not directly applicable to decoder-only architectures. Other approaches apply matrix decomposition techniques such as SVD to hidden states (Chen et al., 2024; Du et al., 2024; Park et al., 2025), isolating subspaces thought to capture hallucination-specific features. These methods assume hallucinations appear in limited proportion within the dataset; if hallucinations dominate or are absent, the identified subspace may fail to characterize them effectively.
Probing-based methods. Recent work probes hidden states directly to capture differences between grounded and hallucinated responses (Oblovatny et al., 2025; Bouchekir et al., 2025; Xu et al., 2023). For instance, Oblovatny et al. (2025) employ MMD to measure distances between prompt and response representations, assuming that grounded responses exhibit larger deviations due to deeper reasoning, whereas hallucinations arise from superficial rephrasing with smaller deviations. However, this assumption is limiting: fabricated responses can also diverge significantly from the prompt, leading to false positives.
2 Methodology
In this section, we describe the proposed method, MMD-Flagger. Given an input document $x$, the starting point of the method is to generate the hypothesis output $y={\rm LLM}(x)$ to be inspected. The next step is to generate additional documents by stochastic sampling at various temperature parameters, which is detailed in Section 2.1. Then, the MMD between $y$ and the newly created groups is computed for each temperature parameter. We detail this computation in Section 2.2. At this point, for a given input / output pair $(x,y)$, we have obtained a series of MMD values corresponding to different temperature settings. We explain our procedure to inspect this trajectory in Section 2.3.
2.1 Decoding strategies and stochastic sampling
In this section, we present the generation process by LLMs considered in this work and detail the role of the decoding strategies.
Text generation with LLMs. We consider LLMs based on the Transformer architecture (Vaswani et al., 2017). Given an input sequence of tokens $x$, the model first processes the sequence through multiple layers of self-attention and feed-forward networks. In encoder–decoder models (Sutskever et al., 2014), the encoder maps $x$ into a contextual representation, which is then attended to by the decoder. In decoder-only models, the same architecture autoregressively encodes and generates tokens within a single stack of Transformer blocks, where causal masking ensures that predictions for each position depend only on preceding tokens.
The decoder generates an outcome sequence autoregressively, conditioned on the prompt and the previously generated tokens. Formally, each token is sampled as
$Y_{t}\sim\Phi\left(Y_{t-1},\ldots,Y_{1},\mathcal{H}(x)\right),$
where $\mathcal{H}(x)$ denotes the model’s hidden representation of the prompt, and $\Phi$ is a stochastic sampler applied to the model’s predictive distribution. In case of the encoder-decoder models, $\mathcal{H}(x)$ is the encoder output.
At each step, the model produces a vector of logits in $\mathbb{R}^{V}$, where $V$ is the vocabulary size. These logits are transformed by a softmax layer into a probability distribution $P\in\mathbb{R}^{V}$ over the next token, from which the sampler $\Phi$ selects the output. This autoregressive process continues until an end-of-sequence (EOS) token is generated.
Stochastic sampling with temperature. The stochastic sampling strategy collects multiple outputs for a given input $x$ that are used to inspect with the hypothesis output $y$. That is, sample a random variable with distribution given by $P$ and select the associated token. We denote $Y_{\text{sto}}=\text{LLM}_{\text{sto}}(x)$ the random output obtained by stochastic sampling.
A variation is introduced by a temperature parameter $\tau>0$. The softmax transformation is modified into $P_{i}=\mathrm{exp}(z_{i}/\tau)/\sum_{j=1}^{V}\mathrm{exp}(z_{j}/\tau)$, where $z_{i}$ represents the logit score for the $i$-th token. As recalled in Section 1, high temperature gives a lot of variability (Yehuda et al., 2024) (equivalent to uniform sampling in the limit), while low temperature degrades to argmax sampling.
Decoding strategies for the hypothesis. We introduce the decoding strategies for selecting the token associated with the highest $P_{i}$ as output. These decoding strategies are used to obtain a token sequence that is inspected if the hallucination exists. We denote this generation process as $y_{\rm hyp}={\rm LLM_{\rm hyp}}(x)$.
A systematic study (Shi et al., 2024) shows that no universally optimal decoding strategy exists; the best choice depends on the model architecture, alignment process, computational resources, and application requirements. As a general guideline, beam search is recommended for aligned models, while stochastic sampling strategies are preferred when computational resources allow, as they often improve diversity and robustness. Beam search is a deterministic decoding algorithm that maintains a set of $w$ candidate sequences, updating them at each time step by retaining the $w$-highest probability continuations (Meister et al., 2020). Although deterministic, beam search and its variants are plausible inspection targets because the token candidates are fundamentally selected based on probability distributions. This probabilistic foundation ensures that MMD-Flagger can inspect hypothesis outputs generated by any decoding method, including deterministic approaches like beam search as well as stochastic sampling strategies such as Top-$k$ sampling (Fan et al., 2018) and nucleus sampling (Top-$p$) (Holtzman et al., 2020).
2.2 Maximum Mean Discrepancy (MMD)
Figure 2: Illustration of the maximum mean discrepancy (MMD). MMD measures the distance between two probability distributions by mapping them into a Reproducing Kernel Hilbert Space (RKHS) and computing the distance.
Maximum Mean Discrepancy (MMD) is a measure of distance between two probability distributions. It relies on a positive semi-definite kernel (Aronszajn, 1950; Hofmann et al., 2008; Smola et al., 2009) to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points (Muandet et al., 2017). A large MMD value indicates significant differences between two distributions. We depict an intuitive presentation of the MMD in Figure 2.
Formal definition. Formally, for a given kernel $k$, the MMD between two distributions $P$ and $Q$ is defined as
[…]
where $A,A^{\prime}\in\mathbb{R}^{D}$ are independent random vectors drawn from $P$, and $B,B^{\prime}\in\mathbb{R}^{D}$ are drawn from $Q$.
Empirical estimate. Given i.i.d. samples ${\bf A}=\{A^{1},\dots,A^{n}\}\sim P$ and ${\bf B}=\{B^{1},\dots,B^{m}\}\sim Q$, one can estimate the MMD by the corresponding empirical averages:
$\begin{split}\widehat{\rm MMD}^{2}_{U}({\bf A},{\bf B})&:=\frac{1}{n(n-1)}\sum_{\begin{subarray}{c}1\leq i,i^{\prime}\leq n\\ i\neq i^{\prime}\end{subarray}}k(A^{i},A^{i^{\prime}})\\ +\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}1\leq j,j^{\prime}\leq m\\ j\neq j^{\prime}\end{subarray}}&k(B^{j},B^{j^{\prime}})-\frac{2}{nm}\sum_{\begin{subarray}{c}i=1,\dots,n\\ j=1,\dots,m\end{subarray}}k(A^{i},B^{j}).\end{split}$ (1)
This is an unbiased estimate of the MMD between two distributions $P$ and $Q$ and converges at the rate $\mathcal{O}_{\mathbf{P}}(\min(n,m)^{-1/2})$ as $n,m\to\infty$ provided that $\sup_{x\in\mathbb{R}^{d}}k(x,x)<\infty$ (Gretton et al., 2012, Theorem 10).
MMD in our setting. Recall that our goal is to measure the similarity between the outputs of $\text{LLM}_{\rm hyp}(x)$ and those of $\text{LLM}_{\rm sto}^{\tau}(x)$. To do this, we consider these two as two probability distributions, and use Eq. (1) to get an estimate of the MMD.
Vector representation of outputs. To capture the rich semantic relationships between generated outputs, we convert the sequences into dense token embedding vectors. The simplest representation is to resort to the word embedding matrix $E\in\mathbb{R}^{V\times D}$ used by LLMs. Thus, each token sequence $y$ is associated with a sequence of embedding vectors
$(e_{1},\ldots,e_{T})=E_{y_{1},:},\ldots,E_{y_{T},:}\in\mathbb{R}^{T\times D}\,.$
To make it a fixed-length vector $h$, we introduce two aggregations of $T$-length vector representations; avg and concat. The avg aggregation is computing the average over $T$ vectors, namely
$h=\frac{1}{T}\sum_{t=1}^{T}e_{t}\in\mathbb{R}^{D}\,.$
The other possibility we investigate is the concat aggregation joins the $T$ embeddings to form
$h=[e_{1};e_{2};\ldots;e_{T};\underbrace{e_{\rm pad};\ldots;e_{\rm pad}}_{T_{\max}-T}]\in\mathbb{R}^{T_{\max}\cdot D}\,$,
where $e_{\rm pad}=\mathbf{0}_{D}$, and $T_{\rm max}$ is given as a hyperparameter. Having the fixed-length vector representation, we denote the set of fixed-length vectors as $H:=\{h_{1},\ldots,h_{n}\}$.
Kernel choice. We briefly introduce the linear and Gaussian kernel. Choices of other kernels are discussed and evaluated in Appendix D.2. The linear kernel (Dot in the following) $k^{\text{lin}}(a,b)=a^{\top}b$ is the simplest choice, and gives rise to simple measures of similarity. Indeed, it is easy to show that the MMD in that case corresponds to the $L^{2}$ distance between $y_{\text{beam}}$ and the barycenter of $Y^{\tau}_{\rm sto}$. On the other side, the Gaussian kernel is often preferred in machine learning applications, notably because of the flexibility of its parameter tuning. For any $a,b\in\mathbb{R}^{D}$, it is defined as
$k^{\text{Gau}}(a,b)=\exp\left(\frac{-\|a-b\|^{2}}{2\gamma^{2}}\right)\,.$ (2)
In the previous display, the bandwidth parameter $\gamma>0$ controls the scaling of input distances before taking the exponential. The median heuristic (Garreau et al., 2017), which sets the length scale to the median of pairwise distances, is commonly used in kernel methods due to its empirical effectiveness. We explored various settings of length scale selections, and selected the best percentile at $25$-th percentile, which is reported in Appendix D.1.
To conclude this section, we note that MMD is a good choice when the data dimensionality is high and the available sample size is limited (Gretton et al., 2012). These MMD’s advantages are well-suited to our setting since the vectorized outputs can have quite a high dimension (especially in the concat setting), while the number of sampled sequences $H^{\tau}_{\text{sto}}$ is strongly limited by our compute budget.
2.3 MMD Flagger
Now, we are equipped with all the machinery required to introduce our proposed method. MMD-Flagger is a simple rule-based approach checking MMD distances $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ for a grid of temperature parameters $T=\{\tau_{1},\dots,\tau\}$. By default, we propose to take $T=\{0.1,\ldots,1.0\}$ with a step of $0.1$. For each $\tau\in T$, we query the model to sample token sequences $Y^{\tau}_{\rm sto}:=\{{y^{\tau}_{\rm sto}}\}_{n=1}^{N}$ and its vector representations $H^{\tau}_{\rm sto}:=\{h_{\rm sto}^{\tau}\}_{n=1}^{N}$ by the stochastic sampling with temperature parameter $\tau$, as described in Section 2.1. We then compute the MMD distances defined by Eq. (1). Namely,
$\begin{split}\mathcal{T}_{\widehat{MMD}}:=(&\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto}),\\ &\ldots,\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto}))\,.\end{split}$
Since an empirical MMD estimate (Eq 1) converges to the true MMD as the sample size increases, the variability in the computed MMD trajectory gradually disappears, as empirically shown in Appendix F.
After computing these values, we detect the inflection point of this trajectory and flag it as a hallucination if it is close to 0. To detect this inflection point, we proceed in the following way: first, we consider a minimal temperature $\tau_{0}$, set by default at $0.11$. Then we compute $\tau_{\text{min}}$ the temperature parameter realizing the minimal MMD distance between the reference and the stochastic samples. In other words, $\tau_{\min}$ is the index of $\mathcal{T}_{\widehat{MMD}}$ with the minimal value. Finally, we flag as hallucination if $\tau_{\min}$ is greater than or equal to $\tau_{0}$. Figure 1 showcases an example of the method at work in both the hallucination and the non-hallucination scenario. We collect the operating procedure in Algorithm 1.
Input: LLM, source input $x$, temperature parameters $T$, number of samples NN, a minimum temperature for detecting the hallucination $\tau_{0}$
Obtain a hypothesis sequence $H_{\text{hyp}}$
foreach $\tau\in T$ do
   Repeat stochastic translation in NN times and obtain $H_{\text{sto}}^{\tau}=\text{LLM}_{\text{sto}}^{\tau}(x)$
   Compute the MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ using Eq. (1)
end foreach
[…]
if $\tau_{\min}>\tau_{0}$ then
 return hallucination
Algorithm 1 MMD-Flagger
3 Empirical Assessment
In this section, we evaluate MMD-Flagger on four hallucination benchmark datasets. All experiments were done on a single NVIDIA Tesla V100 GPU. We report more details on the experimental setup in Appendix A and comparison with SelfCheckGPT (Manakul et al., 2023) leveraging external resources in Appendix C.1.
Table 2: Hallucination detection performance reported as “Recall (Precision).” We mark the best recall in bold. LFAN-HALL, Halomi are machine translation datasets, and ConstFact, FaithBench are summarization datasets.
LFAN-HALL Halomi ConstFact FaithBench
TNG 0.069 (1.000) 0.002 (1.000) - -
Baselines MC-Dsim 0.815 (0.182) 0.553 (0.429) 0.458 (0.750) 0.363 (0.266)
Seq-Logprob 0.850 (0.189) 0.381 (0.273) 0.348 (0.570) 0.363 (0.4)
avg-Gaussian 0.816 (0.129) 0.825 (0.281) 0.904 (0.667) 0.909 (0.555)
Ours avg-Dot 0.803 (0.124) 0.779 (0.278) 0.872 (0.677) 0.909 (0.625)
concat-Gaussian 0.850 (0.134) 0.824 (0.278) 0.879 (0.648) -
concat-Dot 0.833 (0.130) 0.817 (0.282) 0.838 (0.644) -
3.1 Datasets and metrics
We evaluate MMD-Flagger and baselines on four distinct hallucination benchmarks; two for the machine translation task LFAN-HALL (Guerreiro et al., 2023b), Halomi (Dale et al., 2023) and the other two for the summarization task ConstFact (Dreyer et al., 2023), FaithBench (Bao et al., 2025).
LFAN-HALL dataset is a machine translation hallucination detection benchmark. The dataset was generated using an encoder-decoder model provided by the authors, that is a smaller version of a WMT18 translation model with roughly 7777M parameters, produced the $3,413$ translations in the dataset. We note that after our manual review of the annotations, the corrected dataset, included in the supplementary materials, reveals that approximately $8\%$ of the translations are classified as hallucinated.
Halomi dataset is a multilingual hallucination detection benchmark for machine translation, encompassing over $3,000$ sentences across nine languages. The translations were generated using the multilingual encoder-decoder model nllb-200-distilled-600M (Costa-Jussà et al., 2022). In this dataset, approximately $26\%$ of the translations are annotated as hallucinated.
ConstFact dataset evaluates the factual consistency of summaries derived from the CNN/Daily Mail dataset. The summaries within ConstFact were generated using the BART encoder-decoder architecture (Lewis et al., 2020), specifically the bart.large.cnn variant with over 406406M parameters.
FaithBench dataset is a benchmark of evaluating the faithfulness of summaries from decoder-only models. The hypothesis outputs are generated with argmax strategy that does not fit with MMD-Flagger, discussed in Section 4. To tailor this dataset to our needs, we constructed a custom sub-dataset. We randomly selected $25$ source documents and regenerated summaries using the beam search decoding strategy with the Phi-3-mini-4k model (Grattafiori et al., 2024). This is in contrast to the original dataset’s use of the argmax decoding. The model has 3.83.8B parameters. Our newly annotated sub-dataset, with $11$ hallucinated and $14$ faithful summaries, is included in our supplementary materials.
Evaluation Metrics. We use recall and precision as evaluation metrics. They are computed based on the binary labels associated with hallucination / not hallucination given by the datasets. We treat the “hallucination” label as true label. Recall is defined as $\frac{T_{p}}{T_{p}+F_{n}}$, and precision is defined as $\frac{T_{p}}{T_{p}+F_{p}},$ where $T_{p}$ denotes the number of true positives, $F_{p}$ the number of false positives, and $F_{n}$ the number of false negatives.
3.2 MMD-Flagger settings
As temperature parameters, we set $T=\{0.1,0.2,\ldots,0.9,1.0\}$. The set of stochastic sampling sequences $Y_{\text{sto}}^{\tau}$ and corresponding vector representations $H_{\rm sto}^{\tau}$ consist of $25$ examples for each temperature parameter. We obtain the embedding representation from the word embedding layer of the given LLM. The embedding dimensions are $512$ for LFAN-HALL and $1,024$ for Halomi and ConstFact, and $3,072$ for FaithBench. We do not evaluate a vector representation concat of MMD-Flagger on FaithBench due to the computational constraints that the concatenated vector becomes too large.
As described in Section 2.2, we considered the linear (Dot) and Gaussian (Gaussian) kernels in our experiments. To pick the bandwidth $\gamma$ for the Gaussian kernel, we randomly select $200$ (correct) translations as calibration data. The $T_{\rm max}$ of concat is set to the maximum token length among $200$ translations. We conduct experiments with various configurations of choosing the length scale and select the $25$-th percentile of $L^{2}$-distance, instead of the usual median heuristic (Garreau et al., 2017). Supporting experiments for this choice can be found in Appendix D.1. For the Halomi dataset, selecting the calibration data and calculating the length scale is done per source and target language pair.
We also benchmarked MMD-Flagger on internal representations of the model (namely, taking as $h$ vectors of the intermediate hidden layers). We did not find a consistent pattern in our results, while some internal representations occasionally yield better experimental results, see Appendix E.
3.3 Baseline Methods
We considered the baselines proposed by (Guerreiro et al., 2023b); one rule-based method (TNG) and two uncertainty estimation methods (MC-DSim, Seq-Logprob). Let us describe them briefly.
Top n-gram count (TNG(Raunak et al., 2021) is a rule-based method that compares frequencies of $n$-grams between the source and the translation text. TNG flags the translation as the hallucination if certain $n$-gram counts have significant differences. TNG is designed to detect the oscillation type of hallucination; thus, it struggles to flag strong or fully-detached hallucination cases. We set $n=4$ following (Guerreiro et al., 2023b). The evaluation of TNG is omitted on ConstFact and FaithBench datasets as it targets repetition-type hallucinations not present in these datasets.
Dissimilarity of Monte-Carlo hypotheses (MC-DSim(Fomicheva et al., 2020; Zerva et al., 2021) is the average similarity scores between $y_{\text{hyp}}$ and $Y_{\rm MC}$, defined by $\frac{1}{N}\sum_{i=1}^{N}\text{SIM}(y_{\text{hyp}},Y_{\rm MC}),$ where $Y_{\rm MC}$ is a token sequence generated by the beam search using Monte-Carlo Dropout (Gal and Ghahramani, 2016), and $N$ is the number of Monte-Carlo Dropout generations. Following (Guerreiro et al., 2023b), we set $N=10$ and use the METEOR metric (Banerjee and Lavie, 2005) of the nltk package (Bird and Loper, 2004) to compute the similarity score.
Sequence log-probability (Seq-Logprob) is the average of log-probability, defined by $\frac{1}{T}\sum_{k=1}^{T}\log\mathbb{P}\left(y_{k}\middle|y_{<k},x,\theta\right),$ where $x$ is the source text, $y_{k}$ is the $k$-th token in the translation using the beam search, $y_{<k}$ is the previous tokens, $T$ is the length of tokens in the translation, and $\theta$ is the model parameters of the LLM.
MC-DSim and Seq-Logprob are both using a fixed threshold to detect hallucination. As Guerreiro et al. (2023b) suggests, we fix the threshold at the $40$-th percentile among a given dataset. On the LFAN-HALL dataset, the threshold selection is performed using the whole dataset, while we compute the threshold per language pair of the source and target language on the Halomi dataset. The threshold selection is described in more depth in Appendix C.
3.4 Discussion
Experimental results. Table 2 lists the precision and recall scores on the benchmark datasets. Note that the result of MMD-Flagger with concat vector representations is unavailable to FaithBench, as the resulting embedding size becomes prohibitively large and difficult to handle. MMD-Flagger (concat-Gaussian) matches the best baseline in terms of recall on the LFAN-HALL dataset (Seq-Logprob). MMD-Flagger (avg-Gaussian) outperforms baselines in terms of recall and precision on the Halomi, ConstFact, FaithBench datasets. We see these results as very promising, and observe that they are robust with respect to changes in the kernel and aggregation method.
Concatenation versus averaging. When comparing avg and concat, the concat aggregation generally yielded slightly superior or comparable performance to avg. This trend is plausible, as concat retains all individual embeddings, whereas avg aggregates information through a mean operation, potentially losing finer-grained details. Indeed, concat relies on a much larger vector representation of the output documents, of size $D\cdot T_{\text{max}}$. This can become an issue for longer outputs or if the dimension of the word embedding is too high. For instance, in our experiments with the LFAN-HALL dataset, the selected maximum token length ($T_{\text{max}}$) by the calibration data is $105$ tokens. Given that the dimensionality of a word embedding vector $D=512$, the resulting concat vectors have dimension $53,760$. In contrast, the avg aggregation maintains a fixed-length vector size regardless of the token length. Thus, the avg aggregation is generally more practical, even though detection scores can be slightly lower than those of the concat aggregation.
4 Limitations
The first limitation of MMD-Flagger lies in its reliance on the source of hallucination. The method is designed to detect hallucinations by identifying discrepancies between a hypothesis output and stochastic sampling outputs, as reflected in the trajectory of MMD distances. Consequently, if the sampled sequences $Y_{\text{sto}}^{\tau}$ are highly similar to the hypothesis $y_{\text{hyp}}$ for low temperatures, hallucinations can go undetected. It may be the case for an irreducible proportion of hallucinations.
Another limitation is the computational cost of MMD-Flagger, as it requires repeated stochastic sampling across multiple temperature settings. This cost scales with the size of LLM and the length of the generated sequence, making the method less suitable for very large models or long outputs. A practical workaround is to reduce the number of samples or temperature values used when computing the MMD trajectory. However, the stability and reliability of MMD-Flagger under such reduced settings remain an open question. We note that the MMD estimator is characterized by the quadratic complexity; however, the total inference time is overwhelmingly dominated by the cost of LLM output generation, thus rendering the MMD calculation a negligible component of the overall computation.
A limitation of our current framework lies in the deterministic nature of argmax decoding. The MMD-Flagger method is designed to quantify deviations from a stochastic distribution, and thus is not intended to inspect hypothesis outputs generated via argmax selection. The argmax decoding deterministically selects the highest-probability token at each step, producing rigid and often repetitive sequences, in contrast to sampling-based strategies that introduce the necessary variability for our method. Nevertheless, we acknowledge that argmax decoding is still widely adopted in practice for the purposes of reproducibility in benchmarking and evaluation (Bao et al., 2025; Zhou et al., 2025; Chuang et al., 2024).
5 Conclusion
In this work, we introduced MMD-Flagger, a method for hallucination detection based on the Maximum Mean Discrepancy (MMD) metric. The key idea is to compare output sequences generated by a default decoding strategy and stochastic sampling across varying temperature settings. By measuring the distributional distance between these sequences using MMD, MMD-Flagger achieves detection performance on par with the strongest uncertainty-based estimation methods on two machine translation benchmarks. The method’s performance is robust with respect to changes in the output vectorization and choice of the kernel.
Another promising research direction is to investigate the interpretability of MMD-Flagger. One possibility is to leverage recent work in feature importance for MMD-based testing Sutherland et al. (2017); Mitsuzawa et al. (2023). This would allow to pinpoint embedding features responsible for the detection of hallucinated content by MMD-Flagger. Once such variables are located in intermediate layers, interpretability techniques such as Layerwise Relevance Propagation (Montavon et al., 2019) could be applied to analyze how these variables relate to specific encoder layers or segments of the input.
Acknowledgements
This work has been supported by the French government, through the NIMML project (ANR-21-CE23-0005-01). We extend our sincere thanks to Prof. Motonobu Kanagawa for insightful discussions and valuable feedbacks to the conceptualization of this work.
Appendix
This appendix is organized in the following way: Appendix A details the experimental setup. Appendix B describes relabeling procedure of annotation labels of LFAN-HALL dataset. Appendix C contains additional results related to the hallucination detection experiments (Section 3 of the main paper). Appendix H shows qualitative results of the hallucination detection by MMD-Flagger. Comparing kernel functions are described in Appendix D. We show the MMD-Flagger hallucination detection results using vectors from intermediate layers of models in Appendix E. Finally, we investigate the variability of the proposed hallucination detection procedure in Appendix F.
Contents of the Appendix
Appendix A Configurations of experiments
A.1 Computational environment
Experiments are done on the following computational environment: Xeon Platinum 8268 Processor with a single GPU of Tesla V100S-PCIE 3232GB and 755755GB RAM.
A.2 Datasets
LFAN-HALL (Guerreiro et al., 2023b) is a hallucination detection benchmark dataset for the machine translation task from German to English. The dataset consists in $3,413$ machine translations obtained by a transformer-based encoder-decoder model relying on the fairseq package Ott et al. (2019) and beam search, together with associated reference translation and human-annotated hallucination labels. Since we observed some inaccurate annotation labels, we undertook a review of the annotated labels and made corrections. The relabeling process is detailed in Appendix B, and the corrected dataset is contained in the supplementary file code.tar.gz. The dataset defines three types of hallucination labels occupying approximately $8\%$ of all translations, and we unify these three labels as “hallucination.”
Table 3 represents the label distribution in broad categories, while Table 4 shows the label distribution following (Guerreiro et al., 2023b). According to the authors’ hallucination definition, we classify $\{$ error-repetitions, error-strong, error-full }\} into hallucination and $\{$ error named-entities, error-omission }\} into mt-error.
Table 3: Label distribution of the LFAN-HALL dataset. Labels are classified into three categories: no-error, hallucination, and mt-error. The comma “,” represents that the two labels are given.
Label Count Ratio
no-error 2,885 0.8455
hallucination 291 0.0852
mt-error 233 0.0682
hallucination,mt-error 3 0.0008
Table 4: Label distribution of the LFAN-HALL dataset. An example is assigned two labels separated by a command separator.
Label Count Ratio
no-error 2,885 0.8455
error-omission 193 0.0565
error-full 121 0.0354
error-strong 83 0.0243
error-repetitions, error-strong 47 0.0137
error-named-entities 37 0.0108
error-repetitions 29 0.0084
error-full, error-repetitions 11 0.0032
error-named-entities, error-omission 3 0.0008
error-omission, error-strong 1 0.0002
error-omission, error-repetitions 1 0.0002
error full, error-named-entities 1 0.0002
Halomi (Dale et al., 2023) is a hallucination detection benchmark dataset consisting over $3,000$ sentences in $9$ languages. Halomi consists of beam search translations manually annotated for hallucination. The dataset defines three types of hallucinations that occupy approximately $26\%$ of the dataset, and we again unify them as “hallucination.” The label statistics of the Halomi dataset are shown in Table 5.
Table 5: Hallucination label distribution of the Halomi dataset.
Class label Count Ratio
1 No hallucination 2,131 0.74
3 Partial hallucination 291 0.10
2 Small hallucination 225 0.07
4 Full hallucination 218 0.07
ConstFact (Dreyer et al., 2023) is a dataset designed for evaluating the factual consistency of abstractive summarization models. The dataset is derived from the widely-used CNN/Daily Mail (CNN/DM) summarization dataset. Its primary purpose is to assess the trade-off between abstractiveness and factual accuracy in model-generated summaries. The dataset and its code are publicly available in the authors’ repository. [^1] https://github.com/amazon-science/abstractive-factual-tradeoff
The ConstFact dataset consists of 600 human-annotated evaluation records. Each record includes a source document and a model-generated summary. Three human annotators independently evaluated each summary, providing a label that reflects its factual consistency with the original source document. For our evaluation, we define a summary as a hallucination if, at least, one of the three human annotators marked it as “non-factual”. This operational definition allows us to categorize the model-generated summaries as either factually consistent or containing hallucinations, enabling a quantitative analysis of model performance.
FaithBench (Bao et al., 2025) is a benchmark specifically created to evaluate the factual consistency of summaries generated by modern decoder-only models. The dataset provides annotations on the “faithfulness” of model-generated summaries, categorized into four distinct labels: Questionable‘, Benign, Unwanted-Intrinsic, and Unwanted-Extrinsic. This dataset is available in the authors’ public repository. [^2] https://github.com/vectara/FaithBench
For our experiments, we constructed a custom sub-dataset from FaithBench. The original ‘FaithBench‘ dataset’s summaries were generated using a greedy search (argmax) decoding strategy. We regenerated a set of hypothesis summaries using a beam search decoding strategy with the Phi-3-mini-4k model. We randomly selected $25$ source documents from the original FaithBench dataset. Using the same prompts as the original dataset, we generated new summaries for these documents with the Phi-3-mini-4k model. Following the exact annotation scheme and guidelines outlined in the FaithBench, we then labeled the faithfulness of these $25$ newly generated summaries. Our annotation process yielded a total of $11$ summaries containing hallucinations and $14$ that are factually faithful. This new, annotated dataset is included in the supplementary ‘code.tar.gz‘ archive.
A.3 Evaluated models
LFAN-HALL. The translation model is provided by the authors (Guerreiro et al., 2023b). [^3] https://github.com/deep-spin/hallucinations-in-nmt The model is trained by the authors using the fairseq package on the WMT18 DE-EN dataset (Bojar et al., 2018). The model is a Transformer model consisting of hidden vectors of $512$ dimensions, $6$ layers for both the encoder and decoder with $8$ attention heads. The total parameter size of the model is approximately 7777M parameters. The translation speed is roughly $2$ seconds to sample $25$ sequences. In the experiments using vectors from the intermediate layers in Appendix E, the translation speed is significantly slower: approximately $30$ seconds for sampling $25$ sequences. This is because of the necessity of a custom codebase for extracting the vectors from the intermediate layers. Since we, regardless of our best efforts, could not find interfaces of the fairseq package to access and extract the intermediate layers during the inference mode, it is necessary to implement custom code, which causes a significant slowdown.
Halomi. The translation model is the multilingual translation model (Costa-Jussà et al., 2022), and the model is available on the HuggingFace Hub. [^4] https://huggingface.co/facebook/nllb-200-distilled-600M The model is based on the multilingual encoder-decoder model designed for the machine translation task. [^5] The base model in the transformers package is M2M100: https://huggingface.co/docs/transformers/en/model_doc/m2m_100 The model consists of $12$ layers, and each layer generates a vector of $1,024$ dimensions. In this case, it takes approximately $4$ seconds to sample $25$ sequences.
ConstFact. We use BART, an encoder-decoder architecture proposed by Lewis et al. (2020). Following the setup instruction by (Dreyer et al., 2023), we used the bart.large.cnn variant, a large-scale model fine-tuned on the news articles of the CNN/Daily Mail summarization dataset. The model’s parameters total $406,290,432$. The model provides a word embedding vector with a length of $1,024$ for each token. The model file is distributed through the ‘fairseq‘ repository. [^6] https://github.com/facebookresearch/fairseq/blob/main/examples/bart/README.md
Dreyer et al. (2023) introduces a custom decoding method designed to control the abstractiveness of generated summaries. Since ConstFact dataset gives annotation labels on the summaries generated with the control the abstractiveness, we followed the same configuration at generating $y_{\rm hyp}$. Our stochastic sampling of collecting $Y_{\rm sto}^{\tau}$ is not with the control the abstractiveness since authors’ proposed decoding constraints were not designed for these stochastic methods.
FaithBench. The model is Phi-3-mini-4k with 3.8B parameters. This model is available on the HuggingFace Hub. [^7] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct This model provides a word embedding vector in the length of $3,072$ for each token.
A.4 Implementation details
MMD-Flagger. We use an open source codebase to compute the MMD distances (Mitsuzawa et al., 2023). [^8] https://github.com/Kensuke-Mitsuzawa/mmd-two-sample-test-variable-selection The codebase is implemented using PyTorch (Paszke et al., 2017). About the computational speed, MMD-Flagger requires around $1$ second for computing MMD distances of $10$ temperature parameters when the vector aggregation mode is concat and the dimension size is $D=53,760$ and the sample size of $H_{\text{sto}}^{\tau}$ is $25$. Therefore, the required time for computing an MMD distance is around $0.1$ seconds per $\tau$ parameter.
MC-Dsim. We activate the dropout of the model for simulating the stochastic process and obtaining the translation token sequences. For calculating a distance between the token sequences using the drop-out and without, we use the METEOR metric (Banerjee and Lavie, 2005) of the nltk package (Bird and Loper, 2004) to compute the similarity score. [^9] https://www.nltk.org/api/nltk.translate.meteor_score.html
Seq-LogProb. fairseq and transformers packages both directly give the log probability for a generated sequence.
TNG. We use the original implementation of the algorithm provided by the authors (Raunak et al., 2021). [^10] https://github.com/GEM-benchmark/NL-Augmenter/tree/b64a8efebd842848546ec51d3402ba253be12aa7/nlaugmenter/filters/oscillatory_hallucination
Appendix B LFAN-HALL Dataset Modification and Label Correction
Some annotation labels in the LFAN-HALL dataset appear to be inconsistent with the definitions given by the paper. Referring to the annotation guideline provided (Guerreiro et al., 2023b) [^11] https://github.com/deep-spin/hallucinations-in-nmt/tree/main/data, the strongly detached label is applied when there is some content that is supported by the source sentence. However, there is a significant part of the translation that is not (it conveys different meaning to the meaning in the source sentence). Importantly, the content that is hallucinatory does not reflect a genuine effort to translate any part of the source sentence.
In Table 6, we report an example which, in the original datasets, has two labels: omission (of the machine translation error) and strongly detached (of the hallucination). This example is considered to be labeled as the omission (of the machine translation) since the machine translation correctly translated the phrase of the source text A proud celebration, yet the machine translation lacks the phrase that gives cause for celebrating. Therefore, this example is not a case of hallucination.
Table 6: An example of modified annotation labels in LFAN-HALL dataset. The original annotation labels are omission and strongly detached. See the description above.
Source text (German) Reference translation Machine translation
Ein stolzes Jubiläum, das zum Feiern Anlaß gibt. A proud anniversary that gives cause for celebrating. A proud celebration.
We reviewed the annotation labels, focusing on the translations originally labeled as hallucinations, that is, those marked as “strongly detached”, “fully detached”, or “repetition.” Table 7 presents a confusion matrix comparing our verified labels against the original annotation. Notably, substantial discrepancies were found in cases labelled as a combination of “omission” and “strongly detached.” Specifically, 65% (19 examples) of those were re-labelled as “omission,” which is not categorised as hallucination. This may suggest that the annotators have misinterpreted the definition of “strongly detached.” In particular, when a machine translation contains an omission error, the resulting output may appear distant from the source text, potentially leading to misclassification.
Table 7: Confusion matrix comparing verified and original annotation labels in the LFAN-HALL dataset. Cell values indicate the frequency of each annotation label. The symbol “&” denotes the co-occurrence of two error types. Abbreviations are as follows: “NE”: named entity translation error; “omit”: omission error; “rep”: repetition-based hallucination; “full”: fully detached hallucination; “strong”: strongly detached hallucination. Columns are not shown when the frequency is zero. Yellow-highlighted cells denote notable discrepancies between the verified and original annotation labels.
Verified $\backslash$ Orig. full full&NE full&rep NE&strong omit&rep omit&strong rep rep&strong strong
full 117 0 0 0 0 0 0 0 4
full & /NE 0 1 0 0 0 0 0 0 0
full & rep 0 0 11 0 0 0 0 0 0
NE 0 0 0 2 0 0 0 0 1
NE & omit 0 0 0 0 0 0 0 0 1
NE & strong 0 0 0 0 0 0 0 0 0
omit 0 0 0 1 0 19 0 0 1
omit & rep 0 0 0 0 1 0 0 0 0
omit & strong 0 0 0 0 0 1 0 0 0
rep 0 0 0 0 0 0 29 0 0
rep & strong 0 0 0 0 0 0 0 45 2
strong 0 0 0 1 0 8 0 0 69
none 0 0 0 0 0 1 1 0 8
Appendix C Additional experimental results
C.1 Comparison with SelfCheckGPT
We compare the hallucination detection performance of MMD-Flagger against the consistency-based method SelfCheckGPT (Manakul et al., 2023). A key difference is that SelfCheckGPT requires external models (e.g., RoBERTa, T5) for calculating its final score, whereas our MMD-Flagger approach is entirely self-contained, relying only on the internal representations of the generating model.
Setup. We evaluated the BertScore and MQAG variants of SelfCheckGPT, following the authors’ recommended protocol. For a fair comparison, we used the exact same set of $25$ stochastic samples ($\tau=1.0$) across both MMD-Flagger and SelfCheckGPT. To adapt SelfCheckGPT for the binary classification task, we set three distinct thresholds at the $25$-th, $50$-th, and $75$-th percentiles of the generated scores. Due to the high computational cost associated with generating scores for the SelfCheckGPT baseline, all reported metrics were evaluated on a carefully selected data subset; results for MMD-Flagger, MC-DSIM, and Seq-Logprob were correspondingly down-sampled for a direct comparison.
Table 8: The score formats are “Recall (Precision)”. S-GPT is SelfCheckGPT with the threshold value in the parenthesis. MC-Dsim and Seq-Logprob on ConstFact are currently unavailable.
LFAN-HALL Halomi ConstFact
MC-Dsim 0.826 (0.494) 0.612 (0.316)
Seq-Logprob 0.883 (0.481) 0.548 (0.105)
S-GPT BertScore (25) 1.000 (0.281) 1.000 (0.082) 1.000 (0.565)
S-GPT BertScore (50) 1.000 (0.317) 1.000 (0.082) 1.000 (0.575)
S-GPT BertScore (75) 0.948 (0.464) 0.903 (0.084) 0.115 (0.666)
S-GPT MQAG (25) 0.577 (0.312) 0.709 (0.118) 0.536 (0.660)
S-GPT MQAG (50) 0.242 (0.325) 0.322 (0.131) 0.014 (1.000)
S-GPT MQAG (75) 0.051 (0.350) 0.096 (0.176) 0.000 (0.000)
MMD-Flagger (avg-Gaussian) 0.808 (0.362) 0.677 (0.087) 0.855 (0.590)
MMD-Flagger (avg-Dot) 0.845 (0.357) 0.709 (0.092) 0.855 (0.590)
MMD-Flagger (concat-Gaussian) 0.856 (0.370) 0.645 (0.081) 0.884 (0.693)
MMD-Flagger (concat-Dot) 0.852 (0.375) 0.774 (0.100) 0.855 (0.686)
Results Analysis. Detection performance is summarized in Table 8, where scores are reported as Recall (Precision). While SelfCheckGPT frequently achieves a perfect recall of $1.000$ in several low-threshold settings across all three datasets (e.g., S-GPT BertScore (25)), MMD-Flagger overall achieves highly competitive Recall scores to some of SelfCheckGPT’s configurations (MQAG). The core finding is that MMD-Flagger demonstrates strong efficacy, generating comparable detection metrics without the reliance on external scoring models, thereby presenting a more self-contained solution.
C.2 Additional results for the LFAN-HALL Dataset
More intuition on MMD-Flagger. Figure 3 illustrates examples of the intuition of detecting the hallucination by MMD-Flagger.
In detecting the hallucination by MMD-Flagger, the MMD distance at $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is the principal factor of forming the U-shape in the MMD distances trajectory, as shown in Figure 1; the U-shape is formed since $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is larger than $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ values. Figure 3 depicts distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ per ground-truth annotation label. These distributions represent clear differences by the annotation labels; Comparing these distributions, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ with the hallucination label represents the high volume in the center, while the distribution with the correct label depicts the high volume at the quite low $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ value. In case of the machine translation errors, the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ looks quite different from the distribution of the correct label. There may be room for employing MMD-Flagger for detecting the machine translations; However, further studies are required.
Figure 3: Distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ by annotation label. The horizontal axes represent values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$, and the vertical axes are percentages. Comparing two distributions “Correct” and “Hallucination”, we observe frequent larger values of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ in “Hallucination”. This tendency suggests that the U-shape of the trajectory is likely to form at the case of “Hallucination”.
Figure 4 displays the distributions of the $\tau$ values at which the trajectory $\mathcal{T}_{\widehat{MMD}}$ reaches its minimum. In the case of the hallucination label, the minimum MMD values are often observed at $\tau>0.2$, indicating that the $\mathcal{T}_{\widehat{MMD}}$ trajectory tends to exhibit a U-shape, as exemplified in Figure 1.
[…]
Detection Performance per Label.
Table 9 represents the hallucination detection performance per annotation label. The kernel settings of the MMD-Flagger are the same as described in Section 3.2.
Table 9: Detection performance per error label on the LFAN-HALL dataset. The first row represents short names of hallucination labels that represent the detachment levels of the translation text, and the second row is the total count of the annotation label. The hallucination labels are full-d: “full-detachment”, strong-d: “strong-detachment”, rep: “repetition”. The labels of omission and NE (Named Entity) are machine translation errors. The value formats are “true positive / false positive”. We highlight the highest true positive values for each annotation label.
full-d strong-d rep omission NE
Counts 133 131 88 198 41
MMD-Flagger (avg) 115 / 1,739 102 / 1,752 69 / 1,785 117 / 1,737 32 / 1,822
MMD-Flagger (concat) 118 / 1,748 108 / 1,758 69 / 1,797 125 / 1,741 31 / 1,835
MC-DSim 107 / 1,178 110 / 1,175 66 / 1,219 85 / 1,200 18 / 1,267
Seq-Logprob 122 / 1,163 102 / 1,183 63 / 1,222 74 / 1,211 22 / 1,263
TNG 1 / 19 17 / 3 20 / 0 0 / 20 0 / 20
Threshold Selection
Figure 5 represents the uncertainty score distributions and Receiver operating characteristic (ROC) curves of Seq-Logprob and MC-DSim.
(a) Seq-Logprob
(b) MC-DSim
Figure 5: (Top) Score distributions of Seq-Logprob (left) and MC-DSim (right) on LFAN-HALL dataset. The horizontal axis represents an evaluation score for a datapoint, and the vertical axis represents the frequency. The vertical red dashed line is the threshold for the score, selected at $40$-th percentile of the score distribution. (Bottom) ROC curves of Seq-Logprob (left) and MC-DSim (right) on LFAN-HALL dataset. “AUC” in the caption stands for the Area Under the Curve. The orange dashed lines indicate the chance level.
Appendix D Evaluation using various kernel functions
D.1 Comparison of Kernel Length Scale Selection
The “median heuristic” (Garreau et al., 2017) is a widely used method in the kernel literature for selecting the length scale of a kernel function. It computes the pairwise distance matrix (e.g., using the Euclidean distance) over a set of data points and uses the median of these distances as the length scale. While the $50$-th percentile (i.e., the median) is commonly used, other percentiles can also be considered.
In Table 10, we evaluate hallucination detection performance under different configurations of percentile values and length scale shapes (single or vector).
Table 10: Comparison of hallucination detection performance by configuration of a Gaussian kernel function. The vector representation is by the word embedding layer, and vector aggregation is by concatenating vectors of $T$ tokens.
Aggregation Length scale Percentile LFAN-HALL Halomi
concat single 12.5 0.840 (0.132) 0.824 (0.278)
25 0.850 (0.134) 0.824 (0.278)
50 0.837 (0.133) 0.823 (0.279)
62.5 0.830 (0.133) 0.824 (0.280)
75 0.837 (0.134) 0.820 (0.279)
87.5 0.840 (0.136) 0.820 (0.279)
avg single 12.5 0.816 (0.129) 0.826 (0.285)
25 0.816 (0.129) 0.825 (0.281)
50 0.813 (0.130) 0.820 (0.285)
62.5 0.813 (0.130) 0.817 (0.284)
75 0.810 (0.129) 0.814 (0.284)
87.5 0.810 (0.129) 0.809 (0.283)
D.2 Comparison of various kernel functions
We detail the five kernel functions (Schölkopf and Smola, 2002) employed for the MMD estimate, along with the specific hyperparameter ranges explored in our study. In all definitions, $a$ and $b$ denote the input vectors $\mathbb{R}^{D}$.
The Square Exponential Kernel is defined as:
[…]
where $\gamma>0$ is the inverse length-scale parameter. We tested $\gamma\in\{0.001,0.01,0.1,1.0,10\}$.
The Polynomial Kernel is defined as:
$k(a,b;d,\gamma)=\left(\gamma a^{\intercal}b+r\right)^{d},$
where $d$ is the polynomial degree. We tested $d\in\{1,2,3,4\}$ and $r$ (the intercept) $\{0,5,10\}$.
The Laplace Kernel is defined as:
[…]
where $\sigma>0$ is the length-scale parameter. We tested $\sigma\in\{0.1,0.35,0.6,...,2.0\}$.
The Sigmoid Kernel is defined as:
$k(a,b;\alpha,c)=\text{tahn}\left(\alpha a^{\intercal}b+c\right),$
where $\alpha$ is a scaling factor and $c$ is the intercept. We tested $\alpha\in\{0.01,0.1,1.0\}$ and $c\in\{-1.0,0,1.0\}$.
The Matérn kernel (Rasmussen and Williams, 2006) is defined as:
$\begin{split}k_{\nu}(a,b)&=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{\sqrt{2\nu}}{\ell}d(a,b)\right)^{\nu}\\ &K_{\nu}\left(\frac{\sqrt{2\nu}}{\ell}d(a,b)\right),\end{split}$
[…]
$k_{0.5}(r)$ $=\sigma^{2}\exp\left(-\frac{r}{\ell}\right)$
$k_{1.5}(r)$ $=\sigma^{2}\left(1+\frac{\sqrt{3}r}{\ell}\right)\exp\left(-\frac{\sqrt{3}r}{\ell}\right)$
$k_{2.5}(r)$ $=\sigma^{2}\left(1+\frac{\sqrt{5}r}{\ell}+\frac{5r^{2}}{3\ell^{2}}\right)\exp\left(-\frac{\sqrt{5}r}{\ell}\right),$
[…]
Dataset. We use a subset of the Halomi dataset (Dale et al., 2023). The subsets are two translation pairs from eng-Latn to deu-Latn ($146$ examples) and from deu-Latn to eng-Latn ($155$ examples).
The vector representation is by the word embedding layer, and vector aggregation is with the mean over vectors of $T$ tokens.
Evaluation Result. The detection result is reported at Table 11. For each kernel function, we report “recall (precision)” combination regarding the F1 score that is a harmonic mean of precision and recall. We prioritize F1 score over maximizing recall alone, as certain parameter settings achieve perfect recall (1.0) by trivially classifying all examples as hallucinations, which is not practically meaningful.
As shown in Table 11, the Polynomial kernel and the Gaussian kernel achieve the highest recall ($\mathbf{0.827}$). This indicates that a strong capability for detecting hallucinations when parameters are optimally tuned on a labeled calibration dataset. These two kernels require choosing a proper parameter on a calibration dataset having the hallucination label.
The Dot product kernel exhibits slightly inferior performance ($0.793$ recall). However, because the Dot product kernel is non-parametric and requires no calibration data, it stands as a robust and suitable baseline choice when the calibration dataset is unavailable.
Table 11: Comparison of various kernels at Appendix D.2.
Kernel Recall (Precision), parameters
Polynomial 0.827 (0.146), $d=2,r=0$
Gaussian 0.827 (0.138), $25$-th percentile
Dot 0.793 (0.133)
Square Exp. 0.517 (0.091), $\gamma=0.001$
Laplace 0.137 (0.056), $\sigma=0.1$
Matérn 0.241 (0.067), $\nu=0.5$
Sigmoid 0.275 (0.266), $\alpha=0.01,c=-1.0$
Appendix E Hallucination Detection using Vectors from Intermediate Layers
In this section, we report MMD-Flagger detection scores using vectors at intermediate layers of a model. Previous work reports that the model’s intermediate hidden states correlate with the existence of hallucinated token sequences (Dale et al., 2023; Azaria and Mitchell, 2023; Su et al., 2024). Thus, vector representations from the intermediate vector values of the outputs could also contain valuable information for the hallucination detection task.
In Section 2.2, we introduce the vector representation using the word embedding matrix $E\in\mathbb{R}^{V\times D}$. Instead of employing the matrix $E$, we access an intermediate layer of the LLM’s decoder and extract vector representations after the “Add & Norm” operation at each layer, as defined in the Transformer architecture (Vaswani et al., 2017), where “Add & Norm” means the residual connection followed by layer normalization. In this work, the model for the LFAN-HALL dataset has $6$ layers (the dimension size of $512$), and the model for the Halomi dataset has $12$ layers (the dimension size of $1,024$).
We use the Gaussian kernel for the MMD-Flagger. For calculating the length scale, we collect the vector representations of $200$ calibration data and select the $25$-th percentile of $L^{2}$-distance, in the same manner described in Section 3.2.
Table 12 shows the recall and precision scores. The recall scores’ tendencies of two datasets indicate differently. On the LFAN-HALL dataset, the best recall score is at the $1$-th layer, and the recall score gradually becomes lower as the layer goes deeper. On the Halomi dataset, on the contrary, the recall score is low at the earlier layers, nevertheless, the recall score becomes higher as the layer goes deeper. These tendencies may explain a specific layer of the LLM causing the hallucinations; for instance, the first layer of the LFAN-HALL model lets MMD-Flagger plot the U-shape trajectories more often; therefore, the vector representation of the first layer may explain the cause.
Table 12: Comparison of hallucination detection performance by configuration of a Gaussian kernel function. The vector representation is by the word embedding layer, and vector aggregation is by concatenating vectors of $T$ tokens. The symbols “–” represents the absence of corresponding layers in the model.
Layer Number Aggregation Kernel LFAN-HALL Halomi
layer 1 avg Gaussian 0.891 (0.104) 0.676 (0.310)
layer 2 avg Gaussian 0.884 (0.103) 0.673 (0.310)
layer 3 avg Gaussian 0.874 (0.103) 0.681 (0.305)
layer 4 avg Gaussian 0.857 (0.101) 0.685 (0.303)
layer 5 avg Gaussian 0.850 (0.103) 0.694 (0.304)
layer 6 avg Gaussian 0.840 (0.103) 0.697 (0.303)
layer 7 avg Gaussian 0.720 (0.301)
layer 8 avg Gaussian 0.727 (0.296)
layer 9 avg Gaussian 0.736 (0.348)
layer 12 avg Gaussian 0.767 (0.264)
layer 1 concat Gaussian 0.918 (0.104) 0.702 (0.285)
layer 2 concat Gaussian 0.907 (0.101) 0.695 (0.287)
layer 3 concat Gaussian 0.901 (0.100) 0.691 (0.285)
layer 4 concat Gaussian 0.897 (0.100) 0.699 (0.283)
layer 5 concat Gaussian 0.890 (0.099) 0.717 (0.282)
layer 6 concat Gaussian 0.894 (0.101) 0.747 (0.283)
layer 7 concat Gaussian 0.783 (0.282)
layer 8 concat Gaussian 0.816 (0.270)
layer 9 concat Gaussian 0.853 (0.266)
layer 10 concat Gaussian 0.875 (0.262)
layer 11 concat Gaussian 0.895 (0.257)
layer 12 concat Gaussian 0.914 (0.263)
Appendix F Detection Stability of MMD-Hallucination-Flagger
In this section, we discuss the stability of MMD-Flagger in detecting hallucinations, focusing on the impact of randomness and sample size. The detection process involves stochastic sampling, where a random selection of the next token is made based on the probability distribution generated by the softmax function, as detailed in Section 2.1. Our empirical findings suggest that MMD-Flagger is effective in identifying hallucinations with sequences comprising just $25$ tokens.
We used a subset of the LFAN-HALL dataset to iteratively generate token sequences, performing this process 10 times with sample sizes of $\{10,25,50,100\}$. This subset consists of translation examples that all have hallucination labels, totaling $294$ examples; consequently, we only report the recall score.
The vector representation is set to avg, and the kernel employed is a Gaussian kernel. The length scale is determined to be the $25$-th percentile of $200$ calibration data, following the same method described in Section 3.2. The settings for the Gaussian kernel remain consistent across all experiments, ensuring that the computed MMD distances are comparable.
Figure 6 illustrates the bar plots representing recall scores. Notably, the range of recall scores for a sample size of $25$ is closely aligned with that of $100$ samples. This observation suggests that a sample size of $25$ is adequate for effectively detecting hallucinations.
Figure 6: Detection performance of MMD-Flagger using various sample sizes. The horizontal and vertical axes represent the number of samples and the recall score. The bars represent the average of recall scores in $10$ repetitions, and the error bars are the variance.
The trajectory $\mathcal{T}_{\widehat{MMD}}$ consistently exhibits the U-shape, regardless of the sample size. Particularly, the variance, represented by the error bars, decreases with increasing sample size. A trajectory derived from a sample size of $100$ is expected to closely align with the convergence line, indicating a more stable representation of the MMD trajectory.
The MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ plays a significant role in shaping the U-curve of $\mathcal{T}_{\widehat{MMD}}$. Therefore, we analyze the distribution of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ across 294 examples.
Our assumption is that if $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ is close to the converged value, the distance value does not change significantly, regardless of the random seeds used or the sample size. As a result, the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ are expected to exhibit similar shapes.
Figure 7 illustrates the distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ at the first and third repetitions. With a sample size of 10, the two distributions appear quite different; however, at a sample size of 100, the distribution shapes seem to converge.
(a) At 1st repetition using $10$ samples
(b) At 3rd repetition using $10$ samples
(c) At 1st repetition using $25$ samples
(d) At 3rd repetition using $25$ samples
(e) At 1st repetition using $50$ samples
(f) At 3rd repetition using $50$ samples
(g) At 1st repetition using $100$ samples
(h) At 3rd repetition using $100$ samples
Figure 7: Comparisons of distributions of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ over $294$ examples. The left column shows the distributions at first repetition, and the right column is for the one of the third repetition.
Appendix G MMD trajectory smoothing
Table 13: Comparing the hallucination detection of the original MMD trajectory and smoothed trajectories. The score formats are “recall (precision)”.
Vector representation Kernel Smoothing Window size LFAN-HALL
avg dot No smoothing 0.820 (0.125)
dot Average mean 2 0.782 (0.143)
dot Average mean 3 0.734 (0.145)
gaussian No smoothing 0.823 (0.129)
gaussian Average mean 2 0.785 (0.131)
gaussian Average mean 3 0.737 (0.133)
This section introduces the hallucination detection performance after smoothing the MMD trajectory. Figure 8 plots an example of the smoothing, where the blue solid line is the original MMD trajectory, and purple and green dashed lines are smoothed lines. After the smoothing, a temperature $\tau$ at the argmin value change; the red point on the original MMD trajectory, while “x” marks on the smoothed lines. Even applying the smoothing, we do not observe substantial differences, as shown in Table 13.
The MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ can exhibit irregular shapes, as illustrated in Figure 8 and smoothed trajectories by the average window. To address these irregularities, a smoothing method can be applied to create a more uniform trajectory. By utilizing smoothing techniques, the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ becomes easier to manage, which may enhance the ability of MMD-Flagger to detect hallucinations more effectively.
Table 13 compares the hallucination detection performance of the original MMD trajectory and smoothed trajectories. The length scale computation at $25$-th percentile; same as Section 3.2.
Figure 8: An example of the MMD trajectory $\mathcal{T}_{\widehat{MMD}}$ and smoothed trajectories with the average window by $2$ and 33 sizes. The blue solid line is the original MMD trajectory. The purple dashed line is the smoothed line with the window size of 22, and the green dashed line is the smoothed line with the window size of $3$. The red circle is the minimum value of the original MMD trajectory, and black “x” marks are the minimum values on the smoothed trajectory.
Appendix H Qualitative results
In this section, we show examples of detections; the source text and translations are in the left tables, and the corresponding MMD trajectories $\mathcal{T}_{\widehat{MMD}}$ are in the right plots. The examples are all from the LFAN-HALL dataset, and the source language is German. In each table, we show one example of the stochastic sampled translation $y_{\text{sto}}^{\tau=0.1}$.
The detection examples are all with the MMD-Flagger with the vector representation of avg and the kernel setting of gaussian using the $25$-th percentile value for computing the Gaussian kernel length scale, reported in Table 2. We note that the vertical axes in the following Figures are $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H_{\rm sto}^{\tau})$, which represents that we use the beam search for the decoding method at the next token prediction.
Correct hallucination detections (True Positive). Figures 9, 10, 11 illustrate a true positive detection, where the MMD-Flagger successfully identified a hallucinated translation ($y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.1}$) that significantly deviates from the source and reference texts.
Source Nur der Stuck an der Decke uns das Wandabendmahl konnten nicht restauriert werden.
Reference Only the stuccos of the ceiling and the wall could not be restored.
$y_{\text{hyp}}$ The room was very small and the bathroom was very small.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in a quiet area of the city.
Figure 9: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
Source Auf den jeweiligen Dorfplätzen finden sich Alt und Jung zum Schwätzchen und zum Feiern zusammen.
Reference On both places, young and old meet each other to have a chat or to celebrate together.
$y_{\text{hyp}}$ Old and young people will find themselves together in the village’s respective squares for fun and fun.
$y_{\text{sto}}^{\tau=0.1}$ The village’s squares are home to old and young people who want to swim and celebrate.
Figure 10: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
Source Vermutlich überall ist Situation mit Taxi so - wenn man einen Fremdsprachepreis für ein Taxi geht oben spricht.
Reference Probably everywhere situation with taxi is like this - if one speaks a foreign language price for a taxi goes up.
$y_{\text{hyp}}$ The hotel is located in the heart of the city, just a few steps away from the train station.
$y_{\text{sto}}^{\tau=0.1}$ The hotel is located in the heart of the city, just a few steps from the beach and the beach.
Figure 11: True positive Example. The hypothesis output $y_{\rm hyp}$ is hallucinated. Descriptions are found in Appendix H.
Correct translations (True Negative). Figures 12, 13 show true negative cases. As these generated outputs $y_{\rm hyp}$ are correct, the MMD trajectory is the shape of monotonously increasing, and MMD-Flagger does not flag to these hypothesis outputs.
Source Zum Abschluss der Aussprache stellte der Präsident Folgendes fest:
Reference The president concluded the debate as follows:
$y_{\text{hyp}}$ At the end of the debate, the President said:
$y_{\text{sto}}^{\tau=0.1}$ At the end of the debate, the President noted that:
Figure 12: True negative example. The hypothesis output $y_{\rm hyp}$ is correct output. Descriptions are found in Appendix H.
Source Politiken, auf die besonderes Augenmerk zu richten ist, und
Reference policies which should be given special attention, and
$y_{\text{hyp}}$ policies to which particular attention needs to be paid; and
$y_{\text{sto}}^{\tau=0.1}$ policies to which particular attention must be paid; and
Figure 13: True negative example. The hypothesis output $y_{\rm hyp}$ is correct output. Descriptions are found in Appendix H.
Detection by error (False Positive). At Figures 14, 15, the false positive cases are often observed when $y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.1}$ are semantically similar; however, dissimilarity due to different expressions or words. These false-positive errors would be due to the vector representations avg of the word embedding vectors. Due to the aggregation, the Gaussian kernel failed to compute the similarity correctly, and thus, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ becomes a high value.
Source Ich komme zum Schluß noch darauf zu sprechen.
Reference I shall come back to this point again at the end.
$y_{\text{hyp}}$ Finally, I would like to mention this.
$y_{\text{sto}}^{\tau=0.1}$ I will end by mentioning this.
Figure 14: False positive example. The hypothesis output $y_{\rm hyp}$ is correct output, however, MMD-Flagger flags it as the hallucination. Descriptions are found in Appendix H.
Source Wollen sie eine Firma gründen oder umstrukturieren?
Reference Do you intend to market new product or service?
$y_{\text{hyp}}$ Do they want to set up or restructure a company?
$y_{\text{sto}}^{\tau=0.1}$ Do you want to start a company or restructure it?
Figure 15: False positive example. The hypothesis output $y_{\rm hyp}$ is correct output, however, MMD-Flagger flags it as the hallucination. Descriptions are found in Appendix H.
Fail to detect (False Negative). At Figures 16, 17, 18, the false negative errors are often observed when the stochastic sampled translations $y_{\text{sto}}^{\tau=0.1}$ are really similar to the beam search translation $y_{\text{hyp}}$. In this case, $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto})$ becomes a small value and is hard to form a U-shape; therefore, MMD-Flagger failed to detect the hallucination.
Source Februar 1997 angenommenen Entschließung ABl.
Reference 61. adopted on 19 February 1997;
$y_{\text{hyp}}$ Resolution OJ L 208, p.
$y_{\text{sto}}^{\tau=0.1}$ Resolution OJ L 208, p.
Figure 16: False negative example. The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.
Source Oktober zu verlängern.
Reference police and penitentiary.
$y_{\text{hyp}}$ The Commission has decided to extend the programme.
$y_{\text{sto}}^{\tau=0.1}$ The Commission has decided to extend the programme.
Figure 17: False negative example The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.
Source Herr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit
Reference Martin BARTENSTEINFederal Minister for Economic Affairs and Labour
$y_{\text{hyp}}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit.
$y_{\text{sto}}^{\tau=0.1}$ Mr Martin BARTENSTEINBundesminister für Wirtschaft und Arbeit
Figure 18: False negative example. The hypothesis output $y_{\rm hyp}$ is hallucinated, however, MMD-Flagger fails to flag it as the hallucination. Descriptions are found in Appendix H.